### v1

# Excitation Waveforms for Maximum Instantaneous Power Delivery

Abstract
This paper introduces a computational approach to identify performance constraints in the time-domain based on optimizing the excitation waveform. The method builds on an optimization algorithm that has been employed for decades to establish fundamental limits in the frequency domain and this paper showcases its first comprehensive application to time-domain pulses. The method is applied to arbitrarily polarized multiport antennas and arrays. The demonstration performed is based on finding an antenna’s maximum peak radiation intensity in a given direction and time with limited total input energy available. To highlight the generality of the approach, an analysis on finding optimal illumination for antiferromagnetic memory switching is conducted.
Index Terms:
Optimal control, excitation, pulse radiation, peak power, electromagnetic radiation, antenna arrays, magnetic memories.
I Introduction
Short EM (EM) pulse radiation is of significant importance in many fields, including applications involving impulse, short-pulse, or non-sinusoidal radars [1, 2, 3], ground penetrating radars [4] or through-wall imaging radars [5]. In addition to radar applications, an important discipline within the field is EM immunity testing [6]. Moreover, it has played a crucial role in communication systems that seek to increase information transmission rates [1], EM pulse applications, broadband arrays and high-energy directed beam systems [7, 8, 9, 10]. Apart from radioengineering, pulses are used across the whole spectrum of physics including antiferromagnetic switching which is analyzed in this manuscript.
Traditionally designed antennas using continuous wave analysis and synthesis techniques are frequently used to radiate short-duration pulses. Antennas with wide bandwidth are typcially used, relying on the belief that they ensure sufficient performance. A more advanced approach involves matching an antenna excitation signal with its dispersion characteristics. Such an approach yields significantly better results [1, 11, 7]. There are no widely recognized standards for evaluating the performance of an antenna that radiates pulses, although the following are generally accepted: the pulse distortion in the radiation should be minor, the energy of the reflected signal at the input should be small, or the amplitude of the radiated pulse should be large [4].
This paper establishes a computational approach to determine optimality in the radiation of directional pulses, generalizes the methodology, and allows the tool to be used in unconventional scenarios which had not been previously possible. The tool can assess the quality of an arbitrary antenna and give its optimal excitation.
To find the best possible excitation waveform, the optimization framework stems from the development of fundamental bounds on antenna performance in the frequency domain. By including commonly used computation methods, modern approaches, and evaluation techniques such as the EFIE (EFIE[12], MoM (MoM[13], convex optimization [14] and powerful computation units, limitations are often found on arbitrary quantities in various scenarios. A representative example is when efficiency, Q factor, and gain [15, 16] are intrinsically connected to directional pulse radiation which natively requires antennas with sufficient performance in these quantities. The method was initially introduced as antenna current optimization [17], and is based on current density being expressed in terms of EFIE and MoM representing dof (dof) in convex optimization. In state of the art in fundamental bounds in electromagnetism [18], quantity trade-offs [15, 19], tighter bounds including additional constraints [19], or quantities in specific scenarios, such as implanted antennas [20], are evaluated. The phenomenon is also well developed in fields unrelated to antennas, i.e., cloaking [21], heat transfer [22], magnetic trapping [23], and harmonic generation [24].
A few works addressing optimal excitation in the time domain have been done, notably those considering radiated pulse optimization in [8, 25], where up-to-date methodologies are not used. Initial attempts to use modern tools in the optimization of radiated pulses are presented in [26]. The results show optimality in peak radiation intensity in a given direction of linearly polarized dipole-type antennas, such as the dipole, bowtie, and exponentially tapered bowtie. Evaluation consists of a large number of frequency samples computed using EFIE and MoM to optimize the excitation using the  quadratically constrained quadratic program (QCQP[14, 27] in the frequency domain and then mapping by inverse fast Fourier transform (iFFT) to the time domain [26]. The extension of the method to arbitrarily polarized antennas was initialized in [28].
In this paper, in addition to the applicability of the approach to an almost arbitrary problem using any simulation software, multiport antennas and arrays are the novel features described, as discussed in Section IV. The section III presents the framework in detail and follows the flow chart in Fig. 1 where individual steps are given. Most steps are performed using well-developed, proven techniques in the frequency domain for antenna analysis and computation of fundamental bound. The procedure allows for direct control of the frequency range of the optimal excitation pulse which is always band limited, i.e., angular velocity $\omega=2\pi f$ and frequency $f$ are considered from $f_{\mathrm{min}}$ to $f_{\mathrm{max}}$, given by realistic values [^1] For instance, regarding antenna radiation, the frequency band is constrained using step-recovery diodes [29] with $f_{\mathrm{min}}=0\,\mathrm{Hz}$ and $f_{\mathrm{max}}=3.8\,\mathrm{GHz}$. for pulse generators.
antenna S.1 EM solver characte-ristics S.2 interpolation operators $(\omega)$ S.3 projection basis functions S.4 optimization QCQP solution $(\omega)$ S.5 iFT solution $(t)$
Figure 1: Flow chart of the computation steps S.1 – S.5, starting from the given antenna and port with the characteristic impedance of the feed transmission line and finishing with the optimal solution in the time domain, i.e., the optimal excitation and performance value.
The remainder of the paper is organized as follows. Section II describes the history of how to generate excitation for optimal pulses. The established computation framework for optimal directionally radiated pulses is presented, tested, and verified in an example in Section III, whereby a comparison of the performance of different antenna designs allows the best one to be chosen, and assists in deciding on the optimal layout of an array for pulsed radiation. Using the same methodology, the optimal illumination for antiferromagnetic memory switching is found in Section IV. Capabilities, limitations, and possible extensions are discussed in Section V.
II Historical Background
This section recalls work done on optimal pulsed radiation, an oft-ignored topic in recent times, and highlights the novelty provided within this paper.
The critical advancement to uncovering the optimality of antenna pulse radiation lay in using an excitation signal precisely tailored to the specific dispersion characteristics of the antenna [1]. Achieving this would result in a flat frequency response producing the highest possible pulse amplitude and maximizing the strength of the electric field at a given moment and point in the far field. For arbitrarily shaped antennas, with a particular emphasis on planar pulse radiators, the search for optimality is detailed in [1] and provides insight into performance maximization related to size and bandwidth. The methodology assumes an arbitrary planar current distribution confined to the design region. Although this understanding of bounds provides valuable insight, it falls short of synthesizing practical and applicable designs.
In [7], a distinct performance maximum of pulse radiation is initially introduced to examine the optimal input voltage. The metric and computational approach studied remain consistent. This exploits the computation in the frequency domain, using iFFT, and applies variational optimization to maximize the amplitude of the radiated electric field at a specified time and position in the far field with constrained energy [^2] Constraining the root mean square current amplitude over the surface of the radiator was implemented in [1, 7]. and the bandwidth of the applied signal. Compared to [1], the current in [7] is physically realizable bringing this solution closer to practical implementation as long as the feed voltage can be generated or sufficiently approximated. Although the solution in [7] refers to a dipole antenna, it accommodates arbitrary feed point locations, lengths, or impedance loadings. In these works, the current distribution is determined using EFIE and the Galerkin method to employ piecewise sinusoidal expansion [30].
The progression of the study in determining the optimal transient radiation of arbitrary antennas is documented in [8, 31] and extends the analysis to three-dimensional antennas compared to planar radiators in [1]. The objective metric remains the same as before. Furthermore, the study addresses the maximum possible radiated energy density within a specified time interval while considering the effects of introducing constraints on the sidelobes. Physical realizability is emphasized by limiting Q or super-directivity effects. Fundamental limitations are juxtaposed with the performance of optimally fed antennas [7, 31] and dipole arrays [9], all of the same size [8, 31]. This work builds on [32] and employs spherical mode expansion. Limiting the number of spherical mode expansion terms ensures a low antenna Q and mitigates impractical supergain issues. The Chebyshev synthesis procedure is used to constrain the energy radiated in the side lobes.
Supplementary to [7] and in alignment with [8], optimal excitations for dipole arrays aimed at maximizing the amplitude of the transient radiated electric field at a specified time and position in the far field are detailed in [9], and maximizing energy density in a given time interval at a fixed far-field position are detailed in [2]. As with [7], the MoM is employed to calculate the currents on all dipoles, taking into account the effects of mutual coupling. Piecewise sinusoidal expansion modes are used, reflecting the approach in [7]. The relationship between the total radiated field and terminal voltages is established by superposing the far fields of each terminal. This analysis uses a thin-wire approximation for the dipoles and delta gaps to model their feed-voltage sources. The degrees of freedom in this study encompass the magnitudes and phases of the terminal voltages. Furthermore, the investigation is extended to the synthesis of optimal sidelobe patterns.
Alternatively, one might consider whether the radiated electric field at a far zone is equal to a predetermined function of time rather than focusing on maximizing field intensity. This question is addressed in [10]. In specific applications, such a criterion may be more suitable than maximizing the radiated field amplitude or energy density at a particular time and position in the far zone. The solution to this problem is carried out entirely in the time domain by minimizing a particular norm of the difference between the required and actual electric fields at a point lying in the far zone. The paper tackles the inverse source problem, specifically constructing the input voltage for a given far-zone electric field function of time via an iterative optimization algorithm based on the steepest descent.
In addition to finding optimal antenna excitation, there is a search for optimal antenna design with regard to pulse radiation which cannot be accomplished using conventional antenna synthesis techniques [1]. This question is explored in [4] with a focus on the optimality of conical antennas and in [33] with an emphasis on bowtie antennas. The theoretical analysis used to optimize the conical antenna is grounded in the time domain. Geometrical parameters and resistive loading are adjusted to maximize radiated energy for a given excitation. Similarly, dipole shapes are optimized [34] using the conjugate gradient method.
Given that all previous work has focused on radiation in the far-field zone, the study of the near-field zone remains open. Concentrating EM energy in small regions for short intervals has emerging applications in biomedical applications, such as hyperthermia therapy to treat cancer and other diseases [3]. The approach to achieve a high concentration of EM energy in small regions involving free space and non-dispersive materials is presented in [3]. The framework is demonstrated with linear arrays. Intuitively, the timing of the sources must be adjusted to compensate for differences in the propagation path lengths so that transmitted pulses from the array elements are coherently added at the target location.
Moving to communication systems, optimizing the transient waveform should consider both the transmitting and receiving antennas, in contrast to the work mentioned above. In ultra-wideband radio systems, the crucial metric is the maximum amplitude of the received antenna voltage, its sharpness, or energy, as addressed in [35]. Unbeatable target metrics are achieved by finding the best-performing waveform considering the constraints on energy in different stages of the channel. General optimization results are derived for arbitrary antennas, considering the effects of the generator and load impedances using variational methods. The approach is demonstrated with dipoles using MoM [^3] Piecewise sinusoidal basis is used as well, as in [7, 9]. solutions providing the necessary transfer functions and input impedances from the EFIE analysis with the study enriched by closed-form results for short-dipole antennas. Ideally, an ultra-wideband transmitter’s radiated power spectral density should be as flat as possible. However, in reality, ultra-wideband systems are limited by the performance of the antennas which is far from ideal [35].
The precise metric for the sharpness of the received antenna voltage amplitude and its optimization are detailed in [36]. The objective is the maximal derivative of the receiving voltage waveform. A substantial improvement is observed compared to [35] with results validated through measurement.
III Framework
This section develops and verifies the computational approach for optimal excitation and showcases it on pulsed radiation in a chosen far-field direction and time. Various goals and objective functions with multiple constraints can be considered within the approach presented in this article. The practical usage of the method is given in Section IV where several applications are considered.
III-A Methodology
A representative goal used to demonstrate the methodology, which is widely analyzed in this manuscript, is to find an antenna’s maximum peak radiated power in a given direction and time with limited available total input energy available.
Mathematically speaking, the problem reads,
$\begin{split}\max\limits_{\mathbf{a}(t)}\ &r^{2}\boldsymbol{r}_{0}\cdot\boldsymbol{S}(\boldsymbol{r},t_{0}+r/c),\\ \mathrm{s.t.}\ &W_{\mathrm{inc}}=W_{0},\\ &f\in\left[f_{\mathrm{min}},f_{\mathrm{max}}\right],\end{split}$ (1)
where
$\boldsymbol{r}_{0}\cdot\boldsymbol{S}(\boldsymbol{r},t)=\dfrac{1}{Z_{0}}\left|\boldsymbol{E}_{\infty}(\boldsymbol{r},t)\right|^{2}.$ (2)
The incident waves $\mathbf{a}(t)$ on the feeding transmission line connected to the antenna ports are optimized to maximize radial component $\boldsymbol{r}_{0}$ of Poynting vector $\boldsymbol{S}$ at given point $\boldsymbol{r}$ and time $t_{0}+r/c$ which is normalized by the squared distance from source $r^{2}$. For convenience, the observation time is shifted by the free-space propagation delay caused by the speed of light $c$. Impedance $Z_{0}$ in (2) denotes the impedance of the free space and $\boldsymbol{E}_{\infty}$ denotes the electric intensity in the far-field region. Total incident wave energy $W_{\mathrm{inc}}$ is set to be total available energy $W_{0}$. This work considers $50$-Ohm feeding transmission lines and energy $W_{0}=10^{-10}\,\mathrm{J}$.
The solution to such a problem can be achieved by taking the following steps:
  1. S.1
    First, an antenna with defined ports and feeding transmission lines has to be chosen. The setup is analyzed in an arbitrary EM (EM) solver, which can compute the far field in the desired direction for a given frequency and incident waves on the feed lines to the antenna ports. The EM solver provides a set of far-field values. An arbitrary method, capable of evaluating target quantities for a given excitation, can be used, i.e., finite differences in time domain (FDTD) in CST studio suite [37] (CST).
    The electric field in the far-field zone expressed in the frequency domain reads
    $\boldsymbol{E}_{\infty}(\boldsymbol{r},\omega)=\mathbf{H}(\boldsymbol{r},\omega)\,\mathbf{a}(\omega)\mathrm{e}^{-\mathrm{j}\omega\dfrac{r}{c}},$ (3)
    where $\mathbf{H}$ is a far-field transfer matrix obtained from the chosen solver. An example of getting the transfer matrix is the combination of the EFIE and the MoM detailed in Appendix A. Apart from the quantities needed for optimization, antenna properties such as the reflection coefficient or gain at given frequencies can be examined to provide insight into the spectral selectivity of the antenna.
  2. S.2
    The far-field data are interpolated using rational fitting (sparse data) or linear interpolation (dense data) to provide a smooth function of frequency. In this step, the operator for total incident energy is prepared, namely the identity operator in the quadratic form of the incident waves meaning the incident energy is written as a spectral integral
    $W_{\mathrm{inc}}=\dfrac{1}{4\pi}\int\limits_{-\infty}^{\infty}\mathbf{a}^{\mathrm{H}}\left(\omega\right)\mathbf{a}\left(\omega\right)\mathrm{d}\omega.$ (4)
  3. S.3
    The spectral functions representing the necessary operators for optimization, i.e., total incident energy, and the far field are projected onto a set of continuous-band-limited basis functions. The set of basis functions $\xi_{n}(\omega)$ is described in Appendix B. The projection results in expansion coefficients of the basis function grouped in column vector $\mathbf{q}$, where
    $a_{p}(\omega)=\sum\limits_{n}q_{pn}\xi_{n}(\omega).$ (5)
    The vector is constructed in segments with each segment comprising the expansion coefficients of the basis functions for each feed line $p$. Related matrices $\mathbf{W}$ (incident energy) and $\mathbf{S}$ (far-field power) are derived from (4) and (2) together with (3), respectively. See Appendix B for details.
  4. S.4
    The expansion coefficients are optimized to maximize peak radiated power at the given time and direction according to (1) and the equivalent QCQP reads
    $\begin{split}\max\limits_{\mathbf{q}}\ &\mathbf{q}^{\mathrm{H}}\mathbf{S}\mathbf{q},\\ \mathrm{s.t.}\ &\mathbf{q}^{\mathrm{H}}\mathbf{W}\mathbf{q}=W_{0}.\end{split}$ (6)
    The optimization task is resolved using standard methods of the QCQP. For more details, see [38].
  5. S.5
    The solution is found as optimal vector $\mathbf{q}$ in the frequency domain. The linear combination on the feed line then gives the optimal feeding vector which results in maximum peak radiated intensity in the frequency domain using the operators of step S.2.
    The solution is transformed to the time domain via inverse Fourier transform (iFT) either as a single function or by weighted summation images of the basis functions $\xi_{n}(t)$. The far field is also transformed to depict the radiated pulse waveform in the time domain.
III-B Verification
The steps mentioned above are applied to an example of a single dipole antenna. The calculation is simultaneously performed in the Antenna Toolbox for MATLAB [39] (AToM) and CST. The dipole with a length of $15\,\mathrm{mm}$ and a width of $3\,\mathrm{mm}$ is oriented along the $z$ axis. The port is placed in the middle of the dipole and is connected to an ideal $50\,\mathrm{\Omega}$ feed line. Radiation in the spherical coordinate system is studied in the direction $\varphi=0$ and $\theta=\pi/2$.
The antenna including the port is given and all steps visualized in Fig. 1 can be performed. Choosing AToM and CST as the EM solvers and computing the far field for unity excitation at all frequencies results in Fig. 2.
Figure 2: Far field of a dipole antenna with a length of $15\,\mathrm{mm}$ and width of $3\,\mathrm{mm}$ in directions $\varphi=0$ and $\theta=\pi/2$. The antenna is fed through a delta-gap source connected to the feed line, providing a voltage of $1\,\mathrm{V}$ across the entire frequency spectrum.
Since the dipole is aligned along the $z$-axis, only the $\theta$ component of the electric field corresponds to radiation. The curves are in good agreement.
In addition to the far field, an essential known characteristic is total incident energy, represented by the identity operator as seen in (4). This step accomplishes S.1.
Since the frequency data from CST are dense due to the 1001 samples, only linear interpolation is used. Data resulting from AToM are sparse and rational fitting is used for interpolation [^4] Vector fitting [40] could be applied to enforce passivity and causality..
Metrics such as far field and total input power are projected to the band-limited functions defined in Appendix B resulting in matrices as denoted in S.3. The matrix operators allow for optimization S.4 which gives the optimal solution in the frequency domain.
Since the basis functions used for projection are orthogonal and an identity operator represents total input power, the matrix for total input power is an identity matrix. The optimal excitation is a complex conjugate of the far field [1] in Fig. 2. Total input energy is $W_{0}=10^{-10}\,\mathrm{J}$. Meeting the constraint in (6) is ensured by scaling the magnitude of the incident wave. The optimal waveform in the time domain is the result of iFT in S.5, and is plotted in Fig.3.
Figure 3: Optimal waveform of the incident wave gained from analysis in different solvers. In both cases of AToM and CST, optimization is performed in MATLAB.
The agreement of different solvers is more than satisfactory.
The last step is the computation of the radiated pulse, shown in Fig. 4.
Figure 4: Optimal radiated waveform for $t_{0}=0\,\mathrm{s}$ at distance $r$ gained from iFT of the far field for AToM data and monitoring the far field during simulation in CST, where the optimal incident wave is applied as excitation.
The similarity of the data computed by AToM with iFT and a synthetic measurement of the radiated pulse by optimal waveform excitation in CST verifies and, even more, generalizes the method for determining optimal pulses studied for decades [7, 9, 31, 26, 28]. In addition, this framework allows us to find an optimal excitation using an arbitrary solver and do it for complex scenarios, as demonstrated in the following section.
IV Representative Examples
Multiple antennas and arrays are analyzed using the presented approach. Moreover, this section extends the study to antiferromagnetic memory switching where the target is the electric field magnitude at a chosen point excited by a pulsed incident field. Additional research is also conducted on the energy delivered to the given point during the chosen time interval. Since the solution is independent of the solver, as verified in the previous section, the most suitable one can be used for each scenario.
IV-A Far-Field Pulsed Radiation
IV-A1 Singleport Antennas
As has been concluded in previous work, the bowtie dipole antenna exhibits good performance among planar antennas studied for pulsed radiation in the direction orthogonal to the design plane [26, 28].
In this section, the following antennas are considered: a self-grounded bowtie antenna, a conical dipole antenna, a conical spiral antenna, and a logarithmic periodic antenna. The optimal waveform is determined to maximize the radiated peak power. None of the antennas are topologically optimized for the target. This example aims to demonstrate the utility of the approach and compare the best potentially achievable performance of the proposed antennas, including their optimal excitation waveform.
All samples can be circumscribed by the smallest sphere with a radius of $a=36\,\mathrm{mm}$. The allowed frequency band of the input signal is chosen between $0\,\mathrm{GHz}-15\,\mathrm{GHz}$. The antennas are simulated as if made of copper and connected to the $50\,\Omega$ feeding line. The observation point of the pulse radiation is always chosen in a direction native to each antenna.
The self-grounded bowtie is taken as proposed in [41], the model is created and simulated in CST and can be found in Fig. 5.
Figure 5: Self-grounded bowtie antenna with matching circuit as described in [41].
The authors ascribe great performance in pulsed radiation to this format.
A planar version of the bowtie antenna is analyzed in [26, 28] in relation to pulsed radiation. To take advantage of the volume of the circumscribed sphere, a volumetric bowtie dipole antenna is considered and scaled to fit the sphere. See Fig. 6, which is taken from CST, where the simulation is performed.
Figure 6: Volumetric bowtie dipole antenna.
In [42], the conical spiral antenna is proposed for ground-penetrating radars, an ultra-wideband application. The conical spiral radiates a circularly polarized field compared to the other designs which operate in a linearly polarized regime. The antenna proposed in [42] is scaled to fit the circumscribing sphere mentioned above. The simulation is again performed in CST and shown in Fig. 7.
Figure 7: Conical spiral antenna for ground penetrating radar [42].
The periodic logarithmic antenna, recognized as an ultra-wideband antenna [43], is extensively utilized in electromagnetic emission and immunity testing [6]. Therefore, a wire dipole array is a potential candidate for pulsed radiation and is compared with other antennas. The array is designed to fit the circumscribing sphere and exhibits a decay of $0.9$ in element lengths and an opening angle of $15$ degrees. The strip width equals $0.017$ times the length of the longest element and the number of elements is 20. The antenna is simulated using AToM and shown in Fig. 8.
Figure 8: Logarithmic-periodic dipole wire antenna designed according to [43]. The antenna is discretized in the mesh of $1528$ triangles.
The question remains as to which antenna performs better. A comparison is shown in Fig. 9.
Figure 9: Normalized optimal radiated pointing vector at chosen distance $r$ of the compared antennas.
The logarithmic periodic antenna exceeds the other designs in performance. The performance of the self-grounded bowtie antenna is about two times lower and the other designs perform even worse. This finding could prompt efforts to develop improved designs.
Regarding the origins of what constitutes good performance, the realized gain in the given direction accounts for the pulse’s sharpness and energy. The emerging differences of the antennas are depicted in Fig. 10 which justifies the performance differences presented in Fig. 9.
Figure 10: The realized gain of the antenna, which incorporates both the antenna gain and the effects of impedance matching, is a critical factor for achieving optimal performance.
The highest and flattest realized gain gives the best performance which provides some direction for improved future designs.
In contrast, phase properties must be considered once an optimal excitation signal is generated. Abrupt changes in phase require more complex excitation than antennas with linear phase progress which prefer short excitation waveforms. The optimal excitation waveforms shown in Fig. 11 illustrate the ability of the approach to determine the optimal excitation for an arbitrary setup and how realistic it is to achieve the best performance of the given antennas.
Figure 11: Optimal excitation signals achieving the highest performance in pulsed radiation for individual antennas.
The significant difference in excitation signals is visible for the bowtie and periodic antennas. Sharp excitation is preferred by bowtie antennas as compared to periodic antennas which require a longer waveform for pulsed radiation. The delay in the antennas is also noteworthy as the bowtie antenna immediately radiates energy, whereas the grounded bowtie excitation signal shows a delay in the matching circuit. Meanwhile, the periodic antennas, namely the logarithmic and conical spiral ones, accumulate energy first.
The comparison of the antennas and the finding of the optimal excitation using the approach proposed in this paper show that antenna design and pulse generation capabilities are closely related. The information that can be provided by this computational approach can provide essential information in designing antennas that perform well in pulsed radiation.
IV-A2 Multiport Systems
The extension to multiport antennas or arrays is straightforward, as with dipole arrays within this section. The matrix formulation is also shown in Appendices A and B.
The goal of this section is to show whether end-fire or broad-side pulsed radiation perform better in the setup of the four active dipoles discussed in Section III. The dipoles are spaced half a wavelength at $1\,\mathrm{GHz}$. Again, a band limited from above by frequency $3.8\,\mathrm{GHz}$ is considered.
The optimal radiated pulses of all configurations are compared in Fig. 12 where the magnitudes of the radiated Poynting vector peaks $S$ are given and normalized by radial distance $r$. The results show that broadside radiation performs better than end-fire.
Figure 12: Normalized Poynting vector of a single dipole antenna and array radiated in the end-fire and broadside direction.
Figure 13: Optimal excitation signals for end-fire and broadside pulsed radiation.
Considering the same input energy, the end-fire array increases the performance by more than a factor of four as compared to a single dipole. In the case of a broadside array, the performance boost reaches a factor of six. The difference may be due to the coupling of the field with the elements of the arrays which can be deduced from the optimal excitation waveforms in Fig. 13.
The excitations to all ports are almost identical in broadside operation. For the end-fire array, the delay equal to the distance between the dipoles multiplied by the speed of light is observed. The delays agree with the direction of end-fire radiation which points from dipole port 1 to dipole port 4. The magnitude of the excitation pulse is increased from dipole port 1 to dipole ports 2, 3, and 4, which indicates that the last element contributes the most.
IV-B Switching Antiferromagnetic Memory
The approach mentioned in this manuscript can be generalized in an arbitrary application where the input-output relation is known. In this section, the switching of an antiferromagnetic memory using a device illuminated by a pulsed laser is analyzed. The data are taken from [44]. The goal is to reach a peak electric field intensity high enough for memory switching.
The device is electrically large so the full simulation is computationally expensive. The experimental illuminating field is a Gaussian beam from a parabolic mirror with a diameter comparable to the cross-section of the device. However, the simulation in CST is done only for the central part of the device, as seen in Fig. 14, and the plane wave is considered as excitation. Due to the large width of the beam as compared to the target, the plane wave is an excellent approximation and simplifies the calculation.
Figure 14: CST model of the central part of the antiferromagnetic memory [44] with zoomed target middle section.
Compared to the original geometry [44] the motive is made three times smaller. This shifts the frequency behavior to higher bands and results in improved performance of the excitation used.
The metric of interest is electric field intensity in CuMnAs layer $E_{\mathrm{CuMnAs}}$ in the center of the motive shown in Fig. 14. The CuMnAs motive is prepared on a GaAs substrate and most of the CuMnAs alloy is covered by Au, except the central part as described in [44]. Incident field $E_{\mathrm{inc}}$ is the excitation, which is represented by plane wave propagation orthogonally toward the motive, and polarization is aligned with two of the electrodes. The delay from the source to the point of observation in CuMnAs is estimated to be $0.2\,\mathrm{ps}$.
The excitation waveforms are plotted in Fig. 15.
Figure 15: Incident plane waves at the position of the observation: (o) Pulse used in [44]. (a) Optimal excitation for the maximum peak. (b) Optimal excitation for the maximum peak with a constraint on the incident wave to concentrate $90\,\%$ of energy in the $2~{}\mathrm{ps}$ interval. (c) Optimal excitation for the maximum peak with a constraint in the target field to concentrate $90\,\%$ of energy in the $14~{}\mathrm{ps}$ interval. (d) Optimal excitation for maximum peak with both constraints in (b) and (c).
The pulses are normalized to the maximum value of $10^{5}\,\mathrm{V/cm}$. The excited field magnitude in CuMnAs is provided in Fig. 16.
Figure 16: Magnitude of electric field in CuMnAs for excitation pulses (o)–(d) in Fig. 15.
The pulse used in [44] is denoted as (o). Using the optimization framework within this paper, maximum switching performance is determined. The objective function is the magnitude of the electric field intensity in CuMnAs. The excitation is limited in frequency from $625\,\mathrm{GHz}$ to $2.8\,\mathrm{THz}$. The constraint on the frequency band pertains to the frequency range of the excitation pulse (o). Incident field energy is limited and excitation pulses are normalized to have a maximum amplitude of $10^{5}\,\mathrm{V/cm}$ as in [44].
One of the optimal pulses presented corresponds to the maximum magnitude of the electric field in CuMnAs in $t=0$ which is denoted as (a). The optimization problem is identical to (6) other than matrix $\mathbf{S}$ being replaced by $\mathbf{E}^{\mathrm{H}}\mathbf{E}$, where the electric field at the target point is represented as $\boldsymbol{E}_{\mathrm{CuMnAs}}(\omega)=\mathbf{E}(\omega)\mathbf{q}$ in the frequency domain, and $\mathbf{q}$ corresponds to the expansion of incident plan wave $a(\omega)$ into the basis functions of Appendix B, as done in (5). The relation is computed using CST which allows us to compute the electric field at a given point as a function of frequency considering unity excitation. Mapping the set of the given basis functions results in the desired matrix. The matrix representing the total energy of the incident plane wave is derived from (4) as demonstrated in Appendix B, with the only variation being the inclusion of the free space impedance in the denominator.
The maximum amplitude of the electric field in the CuMnAs alloy significantly exceeds that of the pulse used (o). However, the excitation pulse has a long duration and is quite complex compared to (o). Moreover, the resulting electric field in CuMnAs has many peaks and a long duration.
This evidence indicates the necessity of introducing constraints on the input and output waveforms. The incident wave is constrained to have $90\,\%$ energy in the interval $[-1.2,0.8]\,\mathrm{ps}$ taking into account the delay of $t_{\mathrm{d}}=0.2\,\mathrm{ps}$. The constraint is implemented using matrix $\mathbf{W}_{\mathrm{T}}$ of total available energy and energy available in the interval [2]
$W_{\mathrm{T}}=\dfrac{1}{4\pi^{2}Z_{0}}\int\limits_{-\infty}^{\infty}\int\limits_{-\infty}^{\infty}\dfrac{2\sin\left[\left(\omega_{1}-\omega_{2}\right)T\right]}{\omega_{1}-\omega_{2}}\mathrm{e}^{-\mathrm{j}\left(\omega_{1}-\omega_{2}\right)t_{\mathrm{d}}}\\ a^{*}\left(\omega_{2}\right)a\left(\omega_{1}\right)\mathrm{d}\omega_{2}\mathrm{d}\omega_{1},$ (7)
where $T=1\,\mathrm{ps}$ is the half-time interval and the asterisk represents a complex conjugation. In the frequency domain, the sinc function represents the process of time windowing. The matrix obtained is the outcome of the procedure illustrated in Appendix B regarding total input power. The corresponding QCQP reads
$\begin{split}\max\limits_{\mathbf{q}}\ &\mathbf{q}^{\mathrm{H}}\mathbf{E}^{\mathrm{H}}\mathbf{E}\mathbf{q},\\ \mathrm{s.t.}\ &\mathbf{q}^{\mathrm{H}}\mathbf{W}\mathbf{q}=W_{0},\\ &\mathbf{q}^{\mathrm{H}}\left(0.9\mathbf{W}-\mathbf{W}_{\mathrm{T}}\right)\mathbf{q}\leq 0.\end{split}$ (8)
The optimization solution is denoted as (b). The incident field waveform is now squeezed in the given narrow interval and is similar to the original pulse (o). However, the electric field in CuMnAs reaches a higher peak value
The same step with a time limitation can be done with the electric field in CuMnAs as follows
$\begin{split}\max\limits_{\mathbf{q}}\ &\mathbf{q}^{\mathrm{H}}\mathbf{E}^{\mathrm{H}}\mathbf{E}\mathbf{q},\\ \mathrm{s.t.}\ &\mathbf{q}^{\mathrm{H}}\mathbf{W}\mathbf{q}=W_{0},\\ &\mathbf{q}^{\mathrm{H}}\left(0.9\mathbf{U}-\mathbf{U}_{\mathrm{T}}\right)\mathbf{q}\leq 0,\end{split}$ (9)
where $\mathbf{U}$ and $\mathbf{U}_{\mathrm{T}}$ are matrices of the total squared electric field and the squared electric field in the given time interval for $T=7\,\mathrm{ps}$. The matrices are evaluated equivalently as $\mathbf{W}$ and $\mathbf{W}_{\mathrm{T}}$ except $t_{\mathrm{d}}$. The electric field at the target point as a function of frequency and unity excitation is taken from CST. The function is multiplied by its complex conjugate image and projected into the set of basis functions $\xi_{n}$ from Appendix B.
$\mathbf{U}=\dfrac{1}{4\pi}\int\limits_{-\infty}^{\infty}a^{*}\left(\omega\right)\boldsymbol{E}_{\mathrm{CuMnAs}}^{\mathrm{H}}(\omega)\boldsymbol{E}_{\mathrm{CuMnAs}}(\omega)a\left(\omega\right)\mathrm{d}\omega.$ (10)
Matrix $\mathbf{U}_{\mathrm{T}}$ differs only by introducing integration over two distinct angular frequencies $\omega_{1},\omega_{2}$ and the sinc function in the expression as in (7).
The results are presented in (c) which shows the excitation again spanning a long time, whereas the waveform in CuMnAs outside of the prescribed interval is attenuated compared to the amplitude global maxima which are not far from the optimum without the constraint.
The combination of both constraints for excitation in (b) and for the target field in (c) results in another optimization problem whose solution is presented in (d). Since optimization has more constraints than the previous solutions (a), (b), and (c), performance is lower, but still higher than that of the original pulse (o). The excitation mostly spans the narrow interval and the electric field in CuMnAs is also mainly accumulated in the given interval. The disparity between (o) and (d) is minimal, suggesting that enhancement can only be achieved through topological changes.
The current pulse has already reached its design limits. Thus, achieving better performance necessitates either a redesigned system or a less restricted pulse. This showcases the effectiveness of the tool in configuring pulses for antiferromagnetic memory switching and judging topologies for their capabilities.
V Conclusion
A methodology to determine optimal waveforms is presented in this paper with optimized parameters being defined as linear or quadratic terms. The evaluation requires a post-processing step based on convex optimization and the input data can be calculated in any EM simulator.
The tool can serve and provide quantitative results in many scenarios, including pulsed radiation or antiferromagnetic memory switching with pulse lasers, as discussed in this paper, as well as the creation of hotspots in lossy materials using antenna arrays for applications such as hyperthermia therapy in medicine.
The capability of the tool is to find the optimal excitation waveform for an arbitrarily chosen target metric. Moreover, the number of sources is not restricted. In the case of an antenna, matching can be omitted. This leads to fundamental bounds based on current density as known in the frequency domain, but now the current, apart from space, can be controlled in time allowing us to evaluate fundamental bounds in the time domain for a given design region based on the current, although the values would in many cases be overly optimistic.
When the optimized pulsed response is unsatisfactory, the only way to further improve the performance is topology optimization of the radiating body. Topology optimization would change the design according to the results computed by the proposed tool, though such a procedure would be computationally intensive. However, this would open the way to the automated design of pulsed radiators, antiferromagnetic memories, or other devices that rely on pulses in the time domain.
Appendix A Matrix Operators from EFIE and MoM
A representative example of an EM solver capable of providing required antenna characteristics is AToM based on EFIE and MoM. The matrix representation of multiple quantities is presented in [38, Appendix A]. The extraction of essential metrics to the manuscript is presented in this appendix. Multiple ports are considered.
Using EFIE and MoM, the electric field in the far-field zone expressed in the frequency domain reads
$\boldsymbol{E}_{\infty}(\boldsymbol{r},\omega)=\mathbf{F}(\boldsymbol{r},\omega)\mathbf{I}(\omega)\dfrac{\mathrm{e}^{-\mathrm{j}\dfrac{\omega}{c}r}}{r},$ (11)
where $c$ is the speed of light, $\mathbf{F}$ is a far-field matrix, and $\mathbf{I}$ is a column vector of current density expansion coefficients.
The essential quantities are far field as a function of the incident waves from the feeding transmission line and their total available energy in (4). The electric far field is computed from far-field matrix $\mathbf{F}$ and current vector $\mathbf{I}$, see (11). The current is excited by incident waves $\mathbf{a}$ as
$\mathbf{I}\left(\omega\right)=\mathbf{Z}^{-1}\left(\omega\right)\mathbf{P}\mathbf{v}\left(\omega\right),$ (12)
where $\mathbf{Z}$ is the impedance matrix, and $\mathbf{P}$ is the port matrix which assigns port voltage vector values $\mathbf{v}$ to corresponding positions for the matrix multiplication. Port voltages are given by incident waves as
$\mathbf{v}\left(\omega\right)=\mathbf{k}_{\mathrm{i}}^{-1}\left(\omega\right)\mathbf{a}\left(\omega\right),$ (13)
with incident waves $\mathbf{a}$ and $\mathbf{k}_{\mathrm{i}}$ interconnecting the lossless feeding transmission lines of characteristic impedance $Z_{\mathrm{char}}=50\,\Omega$ considered in this manuscript with antenna port voltages [45].
Other antenna characteristics can be computed in the same way. Emerging quantities for this application are antenna gain or port reflection coefficients. The study of the characteristics provides information on the suitability of individual antennas for application.
Appendix B Band Limited Set of Basis Functions
This manuscript uses a set of basis functions for the spectral expansion of antenna characteristics. Complex-valued expansion coefficients are denoted as $q_{n}$. The functions are band-limited, orthogonal, and continuous, and provide well-behaved real-valued images in iFT. Their explicit form reads
$\xi_{n}(\omega)=\dfrac{1}{\sqrt{\omega_{\mathrm{max}}-\omega_{\mathrm{min}}}}\\ \sin{\left(n\pi\dfrac{|\omega|-\omega_{\mathrm{min}}}{\omega_{\mathrm{max}}-\omega_{\mathrm{min}}}\right)}\begin{cases}1,\\ \mathrm{sign}\,(\omega)\mathrm{j},\end{cases}$ (14)
where the upper row corresponds to the real part, while the lower row corresponds to the imaginary part. The values $\omega_{\mathrm{min}}$ and $\omega_{\mathrm{max}}$ define the spectral band interval where the functions are nonzero.
The far-filed operator is projected in the following way for optimized time $t_{\mathrm{opt}}$
$\hat{\mathbf{F}}_{pn}(\boldsymbol{r},\omega)=\\ \int\limits_{-\infty}^{\infty}\mathbf{F}\left(\boldsymbol{r},\omega\right)\mathbf{Z}^{-1}\left(\omega\right)\mathbf{P}\mathbf{k}_{\mathrm{i}}^{-1}\left(\omega\right)\xi_{n}\left(\omega\right)\mathrm{e}^{\mathrm{j}\omega t_{\mathrm{opt}}}\,\mathrm{d}\omega,$ (15)
where $\mathbf{a}_{p}$ excites only the chosen feeding transmission line $p$ by unity amplitude. Here, it is presented for operators from MoM, see Appendix A, but the same methodology can be applied to antenna characteristics from an arbitrary EM solver. The far-field operator projected has two rows corresponding to polarizations and $pn$ columns with a linear index that combines ports and basis functions. The far-filed operator allows us to define the far-field power matrix
$\mathbf{S}(\boldsymbol{r},\omega)=\hat{\mathbf{F}}^{\mathrm{H}}(\boldsymbol{r},\omega)\hat{\mathbf{F}}(\boldsymbol{r},\omega)$ (16)
in (6). Equivalently, the same process is performed with the total available operator (4) in (6) as follows
$\mathbf{W}_{pm,qn}=\dfrac{1}{4\pi}\int\limits_{-\infty}^{\infty}\xi_{m}^{*}\left(\omega\right)\xi_{n}\left(\omega\right)\mathrm{d}\omega,$ (17)
where $pm,qn$ corresponds to the matrix rows and columns according to feeding lines $p,q$ and basis function $m,n$, respectively. Since the ports are separated and the basis functions are orthogonal to each other with respect to the given integral, the identity matrix represents the total incident wave energy matrix $\mathbf{W}$.

### v2

@@
-Excitation Waveforms for Maximum Instantaneous Power Delivery
+Optimization of Excitation Waveforms for Maximum Instantaneous Field Intensity
 Abstract
@@
 The tool can assess the quality of an arbitrary antenna and give its optimal excitation.
+An essential advantage of the proposed method lies in its ability to assess whether a given system configuration can achieve the desired pulsed performance.
+The method provides critical insights into whether optimization efforts can be confined to pulse shaping or whether a redesign of the system geometry is necessary.
+This dual capability offers valuable guidance for both waveform optimization and structural design, facilitating the decision-making process.
+While co-design approaches that simultaneously optimize waveform and antenna design may approach globally optimal solutions, this work focuses specifically on waveform optimization for fixed antenna designs.
+By decoupling waveform optimization from antenna design, the method ensures broad applicability and low computational complexity.
 To find the best possible excitation waveform, the optimization framework stems from the development of fundamental bounds on antenna performance in the frequency domain.
@@
 Capabilities, limitations, and possible extensions are discussed in Section V.
+Although this work primarily focuses on optimizing excitation waveforms to achieve power metrics in far-field scenarios.
+The proposed framework is easily adjustable for near-field energy focusing or similar applications where precise waveform shaping is crucial, given the knowledge of the impulse response in the respective region.
+However, this work does not address the practical aspects of designing pulsed systems but instead provides a versatile theoretical methodology that is applicable in diverse contexts.
 II Historical Background
@@
 This analysis uses a thin-wire approximation for the dipoles and delta gaps to model their feed-voltage sources.
-The degrees of freedom in this study encompass the magnitudes and phases of the terminal voltages.
+The dof in this study encompass the magnitudes and phases of the terminal voltages.
 Furthermore, the investigation is extended to the synthesis of optimal sidelobe patterns.
@@
 Moving to communication systems, optimizing the transient waveform should consider both the transmitting and receiving antennas, in contrast to the work mentioned above.
-In ultra-wideband radio systems, the crucial metric is the maximum amplitude of the received antenna voltage, its sharpness, or energy, as addressed in [35].
+In ultra-wideband (UWB) radio systems, the crucial metric is the maximum amplitude of the received antenna voltage, its sharpness, or energy, as addressed in [35].
 Unbeatable target metrics are achieved by finding the best-performing waveform considering the constraints on energy in different stages of the channel.
@@
 solutions providing the necessary transfer functions and input impedances from the EFIE analysis with the study enriched by closed-form results for short-dipole antennas.
-Ideally, an ultra-wideband transmitter’s radiated power spectral density should be as flat as possible.
-However, in reality, ultra-wideband systems are limited by the performance of the antennas which is far from ideal [35].
+Ideally, an UWB transmitter’s radiated power spectral density should be as flat as possible.
+However, in reality, UWB systems are limited by the performance of the antennas which is far from ideal [35].
 The precise metric for the sharpness of the received antenna voltage amplitude and its optimization are detailed in [36].
@@
 III Framework
-This section develops and verifies the computational approach for optimal excitation and showcases it on pulsed radiation in a chosen far-field direction and time.
+This section develops and verifies the general computational approach for optimal excitation and showcases it on particular example of pulsed radiation in a chosen far-field direction and time.
 Various goals and objective functions with multiple constraints can be considered within the approach presented in this article.
+The essential user input for the computation is the relation between the excitation and the target metric, i.e., a solution to Maxwell’s equations.
+Knowing the system response, almost arbitrary linear scenarios can be studied, including optimization of radiated field intensity, power delivery efficiency, or waveform shaping tailored to specific spatial or temporal requirements.
+This versatility makes the framework applicable to scenarios such as pulse radiation enhancement and electromagnetic field focusing in localized regions.
+In addition to the examples given in the manuscript, notable applications might be the wireless power delivery from the port of an antenna to the port of another antenna and the creation of hot spots in the near field.
 The practical usage of the method is given in Section IV where several applications are considered.
@@
 An arbitrary method, capable of evaluating target quantities for a given excitation, can be used, i.e., finite differences in time domain (FDTD) in CST studio suite [37] (CST).
+This part of the process is computationally the most time-consuming.
+The computational demands grow with wider frequency ranges $f\in\left[f_{\mathrm{min}},f_{\mathrm{max}}\right]$, but these, on the other hand, introduce more dof in the waveform design providing greater flexibility in shaping the radiated pulses.
     The electric field in the far-field zone expressed in the frequency domain reads
-    $\boldsymbol{E}_{\infty}(\boldsymbol{r},\omega)=\mathbf{H}(\boldsymbol{r},\omega)\,\mathbf{a}(\omega)\mathrm{e}^{-\mathrm{j}\omega\dfrac{r}{c}},$ (3)
-    where $\mathbf{H}$ is a far-field transfer matrix obtained from the chosen solver.
+    $\boldsymbol{E}_{\infty}(\boldsymbol{r},\omega)=\mathbf{H}(\boldsymbol{r},\omega)\,\mathbf{a}(\omega)\mathrm{e}^{-\mathrm{j}\omega r/c},$ (3)
+    where $\mathbf{H}$ is a far-field transfer matrix obtained from the chosen solver and $a(\omega)$ represents incident waves in the frequency domain (for example, power waves in the feeding transmission lines), i.e., Fourier transform of $a(t)$.
 An example of getting the transfer matrix is the combination of the EFIE and the MoM detailed in Appendix A.
@@
     The spectral functions representing the necessary operators for optimization, i.e., total incident energy, and the far field are projected onto a set of continuous-band-limited basis functions.
+The projection follows Galerkin method [38, 13].
 The set of basis functions $\xi_{n}(\omega)$ is described in Appendix B.
+For all considered examples, number of basis functions [^4] Later, the number of basis functions is the number of dof in the optimization and determines the time needed for construction $\mathcal{O}(N^{2})$ of matrices and solution $\mathcal{O}(N^{3})$.
+$N\leq 120$ was sufficient.
 The projection results in expansion coefficients of the basis function grouped in column vector $\mathbf{q}$, where
     $a_{p}(\omega)=\sum\limits_{n}q_{pn}\xi_{n}(\omega).$ (5)
-    The vector is constructed in segments with each segment comprising the expansion coefficients of the basis functions for each feed line $p$.
-Related matrices $\mathbf{W}$ (incident energy) and $\mathbf{S}$ (far-field power) are derived from (4) and (2) together with (3), respectively.
-See Appendix B for details.
+    The vector is constructed in segments with each segment comprising the expansion coefficients of the basis functions for each feed line $p$. Related matrices $\mathbf{W}$
+    $W_{\mathrm{inc}}=\mathbf{q}^{\mathrm{H}}\mathbf{W}\mathbf{q}$ (6)
+    (incident energy) and $\mathbf{S}$ (far-field power) are derived via Galerkin method from (4) and (2) together with (3), respectively.
+    As an example, the incident energy (4) is represented by matrix $\mathbf{W}$, the components of which read
+    $\mathbf{W}_{pm,qn}=\dfrac{\delta_{pq}}{4\pi}\int\limits_{-\infty}^{\infty}\xi_{m}^{*}\left(\omega\right)\xi_{n}\left(\omega\right)\mathrm{d}\omega,$ (7)
+    where $pm,qn$ corresponds to the matrix rows and columns according to feeding lines $p,q$ and basis function $m,n$, respectively.
+Kronecker delta $\delta_{pq}$ is zero except when $p$ and $q$ refer to the same port.
+Since the ports are separated and the basis functions are orthogonal to each other with respect to the given integral, the identity matrix represents the total incident energy matrix $\mathbf{W}$.
+Equivalently projction of other operators considered in this methodological example are presented in Appendix C.
   4. S.4
@@
     $\begin{split}\max\limits_{\mathbf{q}}\ &\mathbf{q}^{\mathrm{H}}\mathbf{S}\mathbf{q},\\ \mathrm{s.t.}\ &\mathbf{q}^{\mathrm{H}}\mathbf{W}\mathbf{q}=W_{0}.\end{split}$ (8)
-    The optimization task is resolved using standard methods of the QCQP. For more details, see [38].
+    The optimization task is resolved using standard methods of the QCQP.
+The optimization problem is convex, so the solution is unique [14], which ensures that the maximum is truly achieved [^5] The optimization problem is equivalent to that for determining performance limitations on antenna metrics in the frequency domain [19]..
+Within this paper, the optimization is more than a thousand times faster than the computation of the impulse response of the system.
+For more details about the convex optimization framework, see [39].
   5. S.5
@@
 The far field is also transformed to depict the radiated pulse waveform in the time domain.
+Although described for far-field radiation scenarios, the methodology explained above is general and can easily be applied to different directions and extended to applications such as near-field energy focusing, wireless power transfer, or UWB radars.
+However, we intentionally omit design concerns related to these systems, such as topology optimization of the radiating and receiving antennas.
+We focus solely on feeding waveform optimization, which is, unlike topology optimization, a convex optimization problem.
 III-B Verification
@@
 Radiation in the spherical coordinate system is studied in the direction $\varphi=0$ and $\theta=\pi/2$.
-The antenna including the port is given and all steps visualized in Fig.
+Azimuthal coordinate $\varphi$ references to angle from positive direction of $x$ axis in $xy$ plane and elevation coordinate $\theta$ references to angle from positive direction of $z$ axis.
+The antenna, including the port, is given, and all steps visualized in Fig.
 1 can be performed.
-Choosing AToM and CST as the EM solvers and computing the far field for unity excitation at all frequencies results in Fig.
+Choosing AToM and CST as the EM solvers and computing the far-field component $F_{\theta}$ in elevation direction for unity excitation at all frequencies results in Fig.
 2.
@@
 Figure 6: Volumetric bowtie dipole antenna.
-In [42], the conical spiral antenna is proposed for ground-penetrating radars, an ultra-wideband application.
+In [43], the conical spiral antenna is proposed for ground-penetrating radars, an UWB application.
 The conical spiral radiates a circularly polarized field compared to the other designs which operate in a linearly polarized regime.
@@
 Figure 7: Conical spiral antenna for ground penetrating radar [43].
-The periodic logarithmic antenna, recognized as an ultra-wideband antenna [43], is extensively utilized in electromagnetic emission and immunity testing [6].
+The periodic logarithmic antenna, recognized as an UWB antenna [44], is extensively utilized in electromagnetic emission and immunity testing [6].
 Therefore, a wire dipole array is a potential candidate for pulsed radiation and is compared with other antennas.
@@
 The comparison of the antennas and the finding of the optimal excitation using the approach proposed in this paper show that antenna design and pulse generation capabilities are closely related.
-The information that can be provided by this computational approach can provide essential information in designing antennas that perform well in pulsed radiation.
+The information that can be provided by this computational approach can provide essential insight into designing antennas that perform well in pulsed radiation.
+Notably, within the optimization framework (1), constraints on field intensity at spatial points can be included without compromising generality or convexity.
+Similarly, upcoming sections demonstrate the imposition of energy constraints at the target point within a time window.
+This method can also be extended to additional spatial points outside the main direction and time frames, still using the same framework.
+However, the inclusion of these constraints may result in an increase in the duration of the optimization process.
 IV-A2 Multiport Systems
-The extension to multiport antennas or arrays is straightforward, as with dipole arrays within this section. The matrix formulation is also shown in Appendices A and B.
+The extension to multiport antennas or arrays is straightforward, as with dipole arrays within this section. The matrix formulation is also shown in (7), Appendices A and C.
 The goal of this section is to show whether end-fire or broad-side pulsed radiation perform better in the setup of the four active dipoles discussed in Section III.
@@
 13.
+Although the end-fire and broadside arrays share the same configuration, their characteristics, especially impedance ratios, differ since the dipoles are fed simultaneously.
+In this particular case, the broadside direction performs better than the end-fire orientation.
 The excitations to all ports are almost identical in broadside operation.
@@
 The magnitude of the excitation pulse is increased from dipole port 1 to dipole ports 2, 3, and 4, which indicates that the last element contributes the most.
+The suggested approach can be applied to multi-dimensional arrays or any radiating structures with multiple ports and arbitrary configurations.
+Typically, computational demands grow as the number of elements increases.
+As demonstrated, expanding the number of independent ports results in expressing the excitation waveforms at each port through the basis functions described in Appendix B.
+The expansion coefficients are distinctively grouped into a vector that includes all dof.
 IV-B Switching Antiferromagnetic Memory
@@
 (c) Optimal excitation for the maximum peak with a constraint in the target field to concentrate $90\,\%$ of energy in the $14~{}\mathrm{ps}$ interval.
-(d) Optimal excitation for maximum peak with both constraints in (b) and (c).
+(d) Optimal excitation for the maximum peak with both constraints in (b) and (c).
 The pulses are normalized to the maximum value of $10^{5}\,\mathrm{V/cm}$. The excited field magnitude in CuMnAs is provided in Fig. 16.
-Figure 16: Magnitude of electric field in CuMnAs for excitation pulses (o)–(d) in Fig. 15.
+Figure 16: Magnitude of the electric field in CuMnAs for excitation pulses (o)–(d) in Fig. 15.
 The pulse used in [45] is denoted as (o).
@@
 Mapping the set of the given basis functions results in the desired matrix.
-The matrix representing the total energy of the incident plane wave is derived from (4) as demonstrated in Appendix B, with the only variation being the inclusion of the free space impedance in the denominator.
+The matrix representing the total energy of the incident plane wave is derived from (4) as demonstrated in (7), with the only variation being the inclusion of the free space impedance in the denominator.
 The maximum amplitude of the electric field in the CuMnAs alloy significantly exceeds that of the pulse used (o).
@@
 In the frequency domain, the sinc function represents the process of time windowing.
-The matrix obtained is the outcome of the procedure illustrated in Appendix B regarding total input power.
+The matrix obtained is the outcome of the procedure illustrated in (7) regarding total input power.
 The corresponding QCQP reads
@@
 The combination of both constraints for excitation in (b) and for the target field in (c) results in another optimization problem whose solution is presented in (d).
-Since optimization has more constraints than the previous solutions (a), (b), and (c), performance is lower, but still higher than that of the original pulse (o).
+Since optimization has more constraints than the previous solutions (a), (b), and (c), performance is lower but still higher than that of the original pulse (o).
+That shows that optimization has achieved maximum potential in the given scenario considering the constraints.
 The excitation mostly spans the narrow interval and the electric field in CuMnAs is also mainly accumulated in the given interval.
-The disparity between (o) and (d) is minimal, suggesting that enhancement can only be achieved through topological changes.
+The disparity between (o) and (d) is minimal, so the used design performs almost the best possible.
+This suggests that enhancement can only be achieved through topological changes even more by co-design of the structure and of the pulse.
 The current pulse has already reached its design limits.
@@
 The evaluation requires a post-processing step based on convex optimization and the input data can be calculated in any EM simulator.
-The tool can serve and provide quantitative results in many scenarios, including pulsed radiation or antiferromagnetic memory switching with pulse lasers, as discussed in this paper, as well as the creation of hotspots in lossy materials using antenna arrays for applications such as hyperthermia therapy in medicine.
+The tool can serve and provide quantitative results in many scenarios, including pulsed radiation or antiferromagnetic memory switching with pulse lasers, as discussed in this paper, as well as the creation of hotspots using near-field energy focusing in lossy materials using antenna arrays for applications such as hyperthermia therapy in medicine.
 The capability of the tool is to find the optimal excitation waveform for an arbitrarily chosen target metric.
@@
 This leads to fundamental bounds based on current density as known in the frequency domain, but now the current, apart from space, can be controlled in time allowing us to evaluate fundamental bounds in the time domain for a given design region based on the current, although the values would in many cases be overly optimistic.
-When the optimized pulsed response is unsatisfactory, the only way to further improve the performance is topology optimization of the radiating body.
-Topology optimization would change the design according to the results computed by the proposed tool, though such a procedure would be computationally intensive.
+An important strength of the proposed method lies in its practical utility for evaluating and optimizing radiating systems.
+Given a specific system configuration, the method determines whether the desired pulsed performance can be achieved.
+If the shape and constraints of the system are sufficient, the method identifies the optimal pulse shape to maximize performance.
+Conversely, if the system cannot meet the imposed requirements, the approach highlights these limitations, offering valuable insights into whether a redesign of the system geometry is necessary.
+By decoupling waveform optimization from structural design, the proposed method ensures a globally optimal solution that can be computed efficiently in polynomial time, unlike the iterative and often local solutions associated with topology optimization.
+This characteristic guarantees the general applicability of the method to existing antenna systems without necessitating their redesign, while still ensuring the best achievable performance for a given scenario and set of constraints.
+While this paper focuses on waveform optimization for a fixed transmitting antenna, the methodology has broader potential.
+It can be extended to co-design scenarios in which both antenna and waveform design are considered simultaneously, making it applicable to a wider range of advanced engineering problems.
+In such case, the topology optimization would change the design according to the results computed by the proposed tool, though such a procedure would be computationally intensive.
 However, this would open the way to the automated design of pulsed radiators, antiferromagnetic memories, or other devices that rely on pulses in the time domain.
@@
 A representative example of an EM solver capable of providing required antenna characteristics is AToM based on EFIE and MoM.
-The matrix representation of multiple quantities is presented in [38, Appendix A].
+The matrix representation of multiple quantities is presented in [39, Appendix A].
 The extraction of essential metrics to the manuscript is presented in this appendix.
@@
 Using EFIE and MoM, the electric field in the far-field zone expressed in the frequency domain reads
-$\boldsymbol{E}_{\infty}(\boldsymbol{r},\omega)=\mathbf{F}(\boldsymbol{r},\omega)\mathbf{I}(\omega)\dfrac{\mathrm{e}^{-\mathrm{j}\dfrac{\omega}{c}r}}{r},$ (11)
+$\boldsymbol{E}_{\infty}(\boldsymbol{r},\omega)=\mathbf{F}(\boldsymbol{r},\omega)\mathbf{I}(\omega)\dfrac{\mathrm{e}^{-\mathrm{j}\omega r/c}}{r},$ (13)
 where $c$ is the speed of light, $\mathbf{F}$ is a far-field matrix, and $\mathbf{I}$ is a column vector of current density expansion coefficients.
@@
 The values $\omega_{\mathrm{min}}$ and $\omega_{\mathrm{max}}$ define the spectral band interval where the functions are nonzero.
+Appendix C Projection via Galerkin Method
+Galerkin method [38, 13] converts continuous operators to matrices using the same basis and testing functions.
+The method is applied to all operators using functions introduced in Appendix B.
+The approach is explicitly presented on the far-field operator and the construction of the far-field power matrix.
 The far-filed operator is projected in the following way for optimized time $t_{\mathrm{opt}}$
@@
 $\mathbf{S}(\boldsymbol{r},\omega)=\hat{\mathbf{F}}^{\mathrm{H}}(\boldsymbol{r},\omega)\hat{\mathbf{F}}(\boldsymbol{r},\omega)$ (18)
-in (6). Equivalently, the same process is performed with the total available operator (4) in (6) as follows
-$\mathbf{W}_{pm,qn}=\dfrac{1}{4\pi}\int\limits_{-\infty}^{\infty}\xi_{m}^{*}\left(\omega\right)\xi_{n}\left(\omega\right)\mathrm{d}\omega,$ (17)
-where $pm,qn$ corresponds to the matrix rows and columns according to feeding lines $p,q$ and basis function $m,n$, respectively.
-Since the ports are separated and the basis functions are orthogonal to each other with respect to the given integral, the identity matrix represents the total incident wave energy matrix $\mathbf{W}$.
+in (8).

### v3

@@
-Optimization of Excitation Waveforms for Maximum Instantaneous Field Intensity
+Joint Signal and Topology Optimization for Maximum Instantaneous Field Intensity
 Abstract
-This paper introduces a computational approach to identify performance constraints in the time-domain based on optimizing the excitation waveform.
-The method builds on an optimization algorithm that has been employed for decades to establish fundamental limits in the frequency domain and this paper showcases its first comprehensive application to time-domain pulses.
-The method is applied to arbitrarily polarized multiport antennas and arrays.
-The demonstration performed is based on finding an antenna’s maximum peak radiation intensity in a given direction and time with limited total input energy available.
-To highlight the generality of the approach, an analysis on finding optimal illumination for antiferromagnetic memory switching is conducted.
+This paper introduces a computational approach to identify performance constraints in the time-domain, offering a way to design systems in pulse operation.
+This work presents a comprehensive application of convex optimization to determine fundamental bounds on time-domain waveforms.
+The approach is applied to arbitrarily polarized multiport antennas and arrays, demonstrating their capability in maximizing peak radiation intensity in a specified direction and time under energy constraints.
+This methodology allows us to consider matching, which is crucial in such applications.
+To highlight the generality of the approach, receiving systems are also studied on an example of finding optimal illumination for antiferromagnetic memory switching.
+Thanks to its efficacy, this work enables joint optimization of excitation and system parameters.
 Index Terms:
-Optimal control, excitation, pulse radiation, peak power, electromagnetic radiation, antenna arrays, magnetic memories.
+Co-design, optimal control, excitation, antenna topology, pulse radiation, peak power, magnetic memories.
 I Introduction
@@
 Moreover, it has played a crucial role in communication systems that seek to increase information transmission rates [1], EM pulse applications, broadband arrays and high-energy directed beam systems [7, 8, 9, 10].
-Apart from radioengineering, pulses are used across the whole spectrum of physics including antiferromagnetic switching which is analyzed in this manuscript.
+Apart from radioengineering and far-field operation, EM pulses are used across the whole spectrum of physics, including antiferromagnetic switching [11], which is analyzed in this manuscript.
 Traditionally designed antennas using continuous wave analysis and synthesis techniques are frequently used to radiate short-duration pulses.
-Antennas with wide bandwidth are typcially used, relying on the belief that they ensure sufficient performance.
+Antennas with wide bandwidth are typically used, relying on the belief that they ensure sufficient performance.
 A more advanced approach involves matching an antenna excitation signal with its dispersion characteristics.
@@
 There are no widely recognized standards for evaluating the performance of an antenna that radiates pulses, although the following are generally accepted: the pulse distortion in the radiation should be minor, the energy of the reflected signal at the input should be small, or the amplitude of the radiated pulse should be large [4].
-This paper establishes a computational approach to determine optimality in the radiation of directional pulses, generalizes the methodology, and allows the tool to be used in unconventional scenarios which had not been previously possible.
-The tool can assess the quality of an arbitrary antenna and give its optimal excitation.
+This paper establishes a computational approach to determine optimality in the instantaneous field intensity, generalizes the methodology, and allows the tool to be used in unconventional scenarios which had not been previously possible.
+The tool can assess the quality of an arbitrary pulse system and give its optimal excitation.
+The approach is coupled with system design process to reach better performance.
 An essential advantage of the proposed method lies in its ability to assess whether a given system configuration can achieve the desired pulsed performance.
@@
 This dual capability offers valuable guidance for both waveform optimization and structural design, facilitating the decision-making process.
-While co-design approaches that simultaneously optimize waveform and antenna design may approach globally optimal solutions, this work focuses specifically on waveform optimization for fixed antenna designs.
-By decoupling waveform optimization from antenna design, the method ensures broad applicability and low computational complexity.
-To find the best possible excitation waveform, the optimization framework stems from the development of fundamental bounds on antenna performance in the frequency domain.
-By including commonly used computation methods, modern approaches, and evaluation techniques such as the EFIE (EFIE[12], MoM (MoM[13], convex optimization [14] and powerful computation units, limitations are often found on arbitrary quantities in various scenarios.
-A representative example is when efficiency, Q factor, and gain [15, 16] are intrinsically connected to directional pulse radiation which natively requires antennas with sufficient performance in these quantities.
-The method was initially introduced as antenna current optimization [17], and is based on current density being expressed in terms of EFIE and MoM representing dof (dof) in convex optimization.
-In state of the art in fundamental bounds in electromagnetism [18], quantity trade-offs [15, 19], tighter bounds including additional constraints [19], or quantities in specific scenarios, such as implanted antennas [20], are evaluated.
-The phenomenon is also well developed in fields unrelated to antennas, i.e., cloaking [21], heat transfer [22], magnetic trapping [23], and harmonic generation [24].
 A few works addressing optimal excitation in the time domain have been done, notably those considering radiated pulse optimization in [8, 13], where up-to-date methodologies are not used.
-Initial attempts to use modern tools in the optimization of radiated pulses are presented in [26].
-The results show optimality in peak radiation intensity in a given direction of linearly polarized dipole-type antennas, such as the dipole, bowtie, and exponentially tapered bowtie.
-Evaluation consists of a large number of frequency samples computed using EFIE and MoM to optimize the excitation using the  quadratically constrained quadratic program (QCQP[14, 27] in the frequency domain and then mapping by inverse fast Fourier transform (iFFT) to the time domain [26].
-The extension of the method to arbitrarily polarized antennas was initialized in [28].
-In this paper, in addition to the applicability of the approach to an almost arbitrary problem using any simulation software, multiport antennas and arrays are the novel features described, as discussed in Section IV.
-The section III presents the framework in detail and follows the flow chart in Fig.
-1 where individual steps are given.
-Most steps are performed using well-developed, proven techniques in the frequency domain for antenna analysis and computation of fundamental bound.
-The procedure allows for direct control of the frequency range of the optimal excitation pulse which is always band limited, i.e., angular velocity $\omega=2\pi f$ and frequency $f$ are considered from $f_{\mathrm{min}}$ to $f_{\mathrm{max}}$, given by realistic values [^1] For instance, regarding antenna radiation, the frequency band is constrained using step-recovery diodes [29] with $f_{\mathrm{min}}=0\,\mathrm{Hz}$ and $f_{\mathrm{max}}=3.8\,\mathrm{GHz}$.
-for pulse generators.
-antenna S.1 EM solver characte-ristics S.2 interpolation operators $(\omega)$ S.3 projection basis functions S.4 optimization QCQP solution $(\omega)$ S.5 iFT solution $(t)$
-Figure 1: Flow chart of the computation steps S.1 – S.5, starting from the given antenna and port with the characteristic impedance of the feed transmission line and finishing with the optimal solution in the time domain, i.e., the optimal excitation and performance value.
-The remainder of the paper is organized as follows.
+Initial attempts to use modern tools in the optimization of radiated pulses are presented in [14, 15, 16].
+These efforts, as well as this manuscript, build upon earlier developments of fundamental bounds in the frequency domain [17, 18], which established a foundation for waveform optimization using convex optimization techniques [19].
+This paper demonstrates that the proposed method is adaptable to a variety of problems involving one or more sources using different simulation software.
+It also acts as a basis for automated co-design of pulse excitation and system topology.
+The method and examples in the manuscript are organized as follows.
 Section II describes the history of how to generate excitation for optimal pulses.
-The established computation framework for optimal directionally radiated pulses is presented, tested, and verified in an example in Section III, whereby a comparison of the performance of different antenna designs allows the best one to be chosen, and assists in deciding on the optimal layout of an array for pulsed radiation.
-Using the same methodology, the optimal illumination for antiferromagnetic memory switching is found in Section IV.
-Capabilities, limitations, and possible extensions are discussed in Section V.
-Although this work primarily focuses on optimizing excitation waveforms to achieve power metrics in far-field scenarios.
-The proposed framework is easily adjustable for near-field energy focusing or similar applications where precise waveform shaping is crucial, given the knowledge of the impulse response in the respective region.
-However, this work does not address the practical aspects of designing pulsed systems but instead provides a versatile theoretical methodology that is applicable in diverse contexts.
+The established computation framework is presented in Section III, and tested on an example of optimal directionally radiated pulses in Section IV.
+A comparison of the performance of different antenna designs is shown in Section V.
+Using the same methodology, the optimal illumination for antiferromagnetic memory switching is found later in Section VI illustrating the usage of the methodology for the nearfield and receiving parts of a system.
+The co-design, including pulse and topology optimization, is presented on an antenna array with multiple ports, demonstrating the broad applicability of the approach later in Section VII.
+Capabilities, limitations, and possible extensions are discussed in Section VIII.
 II Historical Background
 This section recalls work done on optimal pulsed radiation, an oft-ignored topic in recent times, and highlights the novelty provided within this paper.
-The critical advancement to uncovering the optimality of antenna pulse radiation lay in using an excitation signal precisely tailored to the specific dispersion characteristics of the antenna [1].
+The critical advancement to uncovering the optimality of antenna pulse radiation lies in using an excitation signal precisely tailored to the specific dispersion characteristics of the antenna [1].
 Achieving this would result in a flat frequency response producing the highest possible pulse amplitude and maximizing the strength of the electric field at a given moment and point in the far field.
@@
 In [7], a distinct performance maximum of pulse radiation is initially introduced to examine the optimal input voltage.
-The metric and computational approach studied remain consistent.
-This exploits the computation in the frequency domain, using iFFT, and applies variational optimization to maximize the amplitude of the radiated electric field at a specified time and position in the far field with constrained energy [^2] Constraining the root mean square current amplitude over the surface of the radiator was implemented in [1, 7].
+The metric and computational approach remain consistent.
+This exploits the computation in the frequency domain, using inverse fast Fourier transform (iFFT), and applies variational optimization to maximize the amplitude of the radiated electric field at a specified time and position in the far field with constrained energy [^1] Constraining the root mean square current amplitude over the surface of the radiator was implemented in [1, 7].
 and the bandwidth of the applied signal.
-Compared to [1], the current in [7] is physically realizable bringing this solution closer to practical implementation as long as the feed voltage can be generated or sufficiently approximated.
+Compared to [1], the current in [7] is physically realizable, bringing this solution closer to practical implementation as long as the feed voltage can be generated or sufficiently approximated.
 Although the solution in [7] refers to a dipole antenna, it accommodates arbitrary feed point locations, lengths, or impedance loadings.
-In these works, the current distribution is determined using EFIE and the Galerkin method to employ piecewise sinusoidal expansion [30].
+In these works, the current distribution is determined using EFIE (EFIE) and the Galerkin method to employ piecewise sinusoidal expansion [20].
 The progression of the study in determining the optimal transient radiation of arbitrary antennas is documented in [8, 21] and extends the analysis to three-dimensional antennas compared to planar radiators in [1].
@@
 Fundamental limitations are juxtaposed with the performance of optimally fed antennas [7, 21] and dipole arrays [9], all of the same size [8, 21].
-This work builds on [32] and employs spherical mode expansion.
-Limiting the number of spherical mode expansion terms ensures a low antenna Q and mitigates impractical supergain issues.
-The Chebyshev synthesis procedure is used to constrain the energy radiated in the side lobes.
 Supplementary to [7] and in alignment with [8], optimal excitations for dipole arrays aimed at maximizing the amplitude of the transient radiated electric field at a specified time and position in the far field are detailed in [9], and maximizing energy density in a given time interval at a fixed far-field position are detailed in [2].
-As with [7], the MoM is employed to calculate the currents on all dipoles, taking into account the effects of mutual coupling.
-Piecewise sinusoidal expansion modes are used, reflecting the approach in [7].
-The relationship between the total radiated field and terminal voltages is established by superposing the far fields of each terminal.
-This analysis uses a thin-wire approximation for the dipoles and delta gaps to model their feed-voltage sources.
-The dof in this study encompass the magnitudes and phases of the terminal voltages.
-Furthermore, the investigation is extended to the synthesis of optimal sidelobe patterns.
-Alternatively, one might consider whether the radiated electric field at a far zone is equal to a predetermined function of time rather than focusing on maximizing field intensity.
-This question is addressed in [10].
-In specific applications, such a criterion may be more suitable than maximizing the radiated field amplitude or energy density at a particular time and position in the far zone.
-The solution to this problem is carried out entirely in the time domain by minimizing a particular norm of the difference between the required and actual electric fields at a point lying in the far zone.
-The paper tackles the inverse source problem, specifically constructing the input voltage for a given far-zone electric field function of time via an iterative optimization algorithm based on the steepest descent.
+As with [7], the MoM (MoM) is employed to calculate the currents on all dipoles, taking into account the effects of mutual coupling.
 In addition to finding optimal antenna excitation, there is a search for optimal antenna design with regard to pulse radiation which cannot be accomplished using conventional antenna synthesis techniques [1].
@@
 Similarly, dipole shapes are optimized [23] using the conjugate gradient method.
-Given that all previous work has focused on radiation in the far-field zone, the study of the near-field zone remains open.
+Given that all previous work has focused on radiation in the far-field zone, the study of the near-field zone remains open and can be addressed by the framework within this manuscript.
 Concentrating EM energy in small regions for short intervals has emerging applications in biomedical applications, such as hyperthermia therapy to treat cancer and other diseases [3].
-The approach to achieve a high concentration of EM energy in small regions involving free space and non-dispersive materials is presented in [3].
-The framework is demonstrated with linear arrays.
-Intuitively, the timing of the sources must be adjusted to compensate for differences in the propagation path lengths so that transmitted pulses from the array elements are coherently added at the target location.
-Moving to communication systems, optimizing the transient waveform should consider both the transmitting and receiving antennas, in contrast to the work mentioned above.
+Moving to communication systems, optimizing the transient waveform should consider both the transmitting and receiving antennas.
+In contrast to the work mentioned above, such an application can be covered by the approach proposed within this manuscript.
 In ultra-wideband (UWB) radio systems, the crucial metric is the maximum amplitude of the received antenna voltage, its sharpness, or energy, as addressed in [15].
@@
 However, in reality, UWB systems are limited by the performance of the antennas which is far from ideal [15].
-The precise metric for the sharpness of the received antenna voltage amplitude and its optimization are detailed in [36].
-The objective is the maximal derivative of the receiving voltage waveform.
-A substantial improvement is observed compared to [35] with results validated through measurement.
-III Framework
-This section develops and verifies the general computational approach for optimal excitation and showcases it on particular example of pulsed radiation in a chosen far-field direction and time.
+To achieve higher performance, antenna–pulse co-design is demanded.
+III Waveform Optimization
+This section develops the general computational approach to maximize the instantaneous field intensity.
 Various goals and objective functions with multiple constraints can be considered within the approach presented in this article.
 The essential user input for the computation is the relation between the excitation and the target metric, i.e., a solution to Maxwell’s equations.
-Knowing the system response, almost arbitrary linear scenarios can be studied, including optimization of radiated field intensity, power delivery efficiency, or waveform shaping tailored to specific spatial or temporal requirements.
-This versatility makes the framework applicable to scenarios such as pulse radiation enhancement and electromagnetic field focusing in localized regions.
-In addition to the examples given in the manuscript, notable applications might be the wireless power delivery from the port of an antenna to the port of another antenna and the creation of hot spots in the near field.
-The practical usage of the method is given in Section IV where several applications are considered.
-III-A Methodology
-A representative goal used to demonstrate the methodology, which is widely analyzed in this manuscript, is to find an antenna’s maximum peak radiated power in a given direction and time with limited available total input energy available.
+Arbitrary linear scenarios can be studied, including optimization of radiated field intensity, power delivery efficiency, or waveform shaping tailored to specific spatial or temporal requirements.
+The methodology developed in this Section ensures that globally optimal time domain excitation is found for a given system.
+The selected objective function $A(\mathbf{a})$ should be either linear or quadratic in the dof (dof), such as instantaneous power flux at given point, and is optimized under a series of constraints, $B_{i}(\mathbf{a})$, that are also quadratic or linear, notable examples being incident energy, delivered energy at a specific time frame, and field values at particular times and positions.
+The dof are represented by incident power waves [24] collected in vector $\mathbf{a}$.
+In contrast to previous work, this methodology allows us to consider matching, which, as shown in subsequent sections, is essential for high performance in the pulse regime.
 Mathematically speaking, the problem reads,
-$\begin{split}\max\limits_{\mathbf{a}(t)}\ &r^{2}\boldsymbol{r}_{0}\cdot\boldsymbol{S}(\boldsymbol{r},t_{0}+r/c),\\ \mathrm{s.t.}\ &W_{\mathrm{inc}}=W_{0},\\ &f\in\left[f_{\mathrm{min}},f_{\mathrm{max}}\right],\end{split}$ (1)
-where
-$\boldsymbol{r}_{0}\cdot\boldsymbol{S}(\boldsymbol{r},t)=\dfrac{1}{Z_{0}}\left|\boldsymbol{E}_{\infty}(\boldsymbol{r},t)\right|^{2}.$ (2)
-The incident waves $\mathbf{a}(t)$ on the feeding transmission line connected to the antenna ports are optimized to maximize radial component $\boldsymbol{r}_{0}$ of Poynting vector $\boldsymbol{S}$ at given point $\boldsymbol{r}$ and time $t_{0}+r/c$ which is normalized by the squared distance from source $r^{2}$.
-For convenience, the observation time is shifted by the free-space propagation delay caused by the speed of light $c$.
-Impedance $Z_{0}$ in (2) denotes the impedance of the free space and $\boldsymbol{E}_{\infty}$ denotes the electric intensity in the far-field region.
-Total incident wave energy $W_{\mathrm{inc}}$ is set to be total available energy $W_{0}$.
-This work considers $50$-Ohm feeding transmission lines and energy $W_{0}=10^{-10}\,\mathrm{J}$.
-The solution to such a problem can be achieved by taking the following steps:
+$\begin{split}\max\limits_{\mathbf{a}\left(t\right)}\ &A(\mathbf{a}),\\ \mathrm{s.t.}\ &B_{i}(\mathbf{a})\leq 0,\forall i\\ &\omega\in\left[\omega_{\mathrm{min}},\omega_{\mathrm{max}}\right].\end{split}$ (1)
+The pulse optimization problem is addressed in the Fourier domain for $\mathbf{a}\left(\omega\right)$ by employing established techniques of convex optimization.
+The solution is further constrained by the available frequency band determined by the specific technology used in the design.
+Specifically, the angular frequency $\omega=2\pi f$ and the frequency $f$ are considered within the range from $\omega_{\mathrm{min}}$ to $\omega_{\mathrm{max}}$, based on practical values [^3] For example, in antenna radiation applications, the frequency band is limited using step-recovery diodes [25], where $f_{\mathrm{min}}=0\,\mathrm{Hz}$ and $f_{\mathrm{max}}=3.8\,\mathrm{GHz}$.
+for pulse generators.
+The problem can be tackled by taking and performing the following steps which are also depicted in the flow chart in Fig. 1.
+system S.1 EM solver characte-ristics S.2 interpolation operators $(\omega)$ S.3 projection basis functions S.4 optimization QCQP solution $(\omega)$ S.5 iFT solution $(t)$ S.6 redesign
+Figure 1: Flow chart of the computation steps S.1 – S.6, starting from the given antenna and port with the characteristic impedance of the feed transmission line and finishing with the optimal solution in the time domain, i.e., the optimal excitation and performance value which may show a need of redesigning the system.
   1. S.1
-    First, an antenna with defined ports and feeding transmission lines has to be chosen.
-The setup is analyzed in an arbitrary EM (EM) solver, which can compute the far field in the desired direction for a given frequency and incident waves on the feed lines to the antenna ports.
-The EM solver provides a set of far-field values.
+    First, a system with defined excitation forms and input channels must be selected.
+The setup is then analyzed using an arbitrary EM solver capable of computing the resulting field intensity at a specified location for a given frequency and a set of incident power waves.
+The solver provides a set of field values corresponding to different excitations.
 An arbitrary method, capable of evaluating target quantities for a given excitation, can be used, i.e., finite differences in time domain (FDTD) in CST studio suite [26] (CST).
 This part of the process is computationally the most time-consuming.
-The computational demands grow with wider frequency ranges $f\in\left[f_{\mathrm{min}},f_{\mathrm{max}}\right]$, but these, on the other hand, introduce more dof in the waveform design providing greater flexibility in shaping the radiated pulses.
-    The electric field in the far-field zone expressed in the frequency domain reads
-    $\boldsymbol{E}_{\infty}(\boldsymbol{r},\omega)=\mathbf{H}(\boldsymbol{r},\omega)\,\mathbf{a}(\omega)\mathrm{e}^{-\mathrm{j}\omega r/c},$ (3)
-    where $\mathbf{H}$ is a far-field transfer matrix obtained from the chosen solver and $a(\omega)$ represents incident waves in the frequency domain (for example, power waves in the feeding transmission lines), i.e., Fourier transform of $a(t)$.
+The computational demands grow with wider frequency ranges $\omega\in\left[\omega_{\mathrm{min}},\omega_{\mathrm{max}}\right]$, but these, on the other hand, introduce more dof in the waveform design providing greater flexibility in shaping the pulses.
+    The provided field vector [^4] It is not strictly required to focus on fields; instead, parameters such as voltage at a port, outgoing power waves along a transmission line, or other related metrics can be analyzed.
+The crucial aspect is the linear relationship between these metrics and incident waves, as described by the transfer matrix.
+expressed in the frequency domain reads
+    $\boldsymbol{E}(\boldsymbol{r},\omega)=\mathbf{H}(\boldsymbol{r},\omega)\,\mathbf{a}(\omega),$ (2)
+    where $\mathbf{H}$ is a transfer matrix obtained from the chosen solver.
 An example of getting the transfer matrix is the combination of the EFIE and the MoM detailed in Appendix A.
-Apart from the quantities needed for optimization, antenna properties such as the reflection coefficient or gain at given frequencies can be examined to provide insight into the spectral selectivity of the antenna.
+Apart from the quantities needed for optimization, system properties can be examined to provide insight into the spectral selectivity of the analyzed system.
   2. S.2
-    The far-field data are interpolated using rational fitting (sparse data) or linear interpolation (dense data) to provide a smooth function of frequency.
-In this step, the operator for total incident energy is prepared, namely the identity operator in the quadratic form of the incident waves meaning the incident energy is written as a spectral integral
-    $W_{\mathrm{inc}}=\dfrac{1}{4\pi}\int\limits_{-\infty}^{\infty}\mathbf{a}^{\mathrm{H}}\left(\omega\right)\mathbf{a}\left(\omega\right)\mathrm{d}\omega.$ (4)
+    The field data are interpolated using rational fitting (sparse data) or linear interpolation (dense data) to provide a smooth function of frequency.
+In this step, the operator for energy and field is prepared, the operators are listed in Appendix B.
   3. S.3
-    The spectral functions representing the necessary operators for optimization, i.e., total incident energy, and the far field are projected onto a set of continuous-band-limited basis functions.
+    The spectral functions representing the necessary operators for optimization, included in the objective function and constraint are projected onto a set of continuous-band-limited basis functions.
 The projection follows Galerkin method [27, 28].
-The set of basis functions $\xi_{n}(\omega)$ is described in Appendix B.
+The set of basis functions $\xi_{n}(\omega)$ is described in Appendix C.
 For all considered examples, number of basis functions [^5] Later, the number of basis functions is the number of dof in the optimization and determines the time needed for construction $\mathcal{O}(N^{2})$ of matrices and solution $\mathcal{O}(N^{3})$.
-$N\leq 120$ was sufficient.
-The projection results in expansion coefficients of the basis function grouped in column vector $\mathbf{q}$, where
+$N\leq 120$ was sufficient and utilizing additional basis functions does not change the result.
+The projection results in expansion coefficients collected in column vector $\mathbf{q}$, where
     $a_{p}(\omega)=\sum\limits_{n}q_{pn}\xi_{n}(\omega).$ (3)
-    The vector is constructed in segments with each segment comprising the expansion coefficients of the basis functions for each feed line $p$. Related matrices $\mathbf{W}$
-    $W_{\mathrm{inc}}=\mathbf{q}^{\mathrm{H}}\mathbf{W}\mathbf{q}$ (6)
-    (incident energy) and $\mathbf{S}$ (far-field power) are derived via Galerkin method from (4) and (2) together with (3), respectively.
-    As an example, the incident energy (4) is represented by matrix $\mathbf{W}$, the components of which read
-    $\mathbf{W}_{pm,qn}=\dfrac{\delta_{pq}}{4\pi}\int\limits_{-\infty}^{\infty}\xi_{m}^{*}\left(\omega\right)\xi_{n}\left(\omega\right)\mathrm{d}\omega,$ (7)
-    where $pm,qn$ corresponds to the matrix rows and columns according to feeding lines $p,q$ and basis function $m,n$, respectively.
-Kronecker delta $\delta_{pq}$ is zero except when $p$ and $q$ refer to the same port.
-Since the ports are separated and the basis functions are orthogonal to each other with respect to the given integral, the identity matrix represents the total incident energy matrix $\mathbf{W}$.
-Equivalently projction of other operators considered in this methodological example are presented in Appendix C.
+    The vector is constructed by blocks, each block representing one source.
+This allows to express functionals from Appendix B as quadratic forms in vectors $\mathbf{q}$.
+Taking energy as an example, the quadratic form reads
+    $W(\boldsymbol{r},T)=\mathbf{q}^{\mathrm{H}}\mathbf{W}(\boldsymbol{r},T)\mathbf{q}$ (4)
+    with matrix elements $\mathbf{W}_{pm,qn}$ where $pm,sn$ corresponds to the matrix rows and columns according to sources $p,q$ and basis function $m,n$, respectively.
   4. S.4
-    The expansion coefficients are optimized to maximize peak radiated power at the given time and direction according to (1) and the equivalent QCQP reads
-    $\begin{split}\max\limits_{\mathbf{q}}\ &\mathbf{q}^{\mathrm{H}}\mathbf{S}\mathbf{q},\\ \mathrm{s.t.}\ &\mathbf{q}^{\mathrm{H}}\mathbf{W}\mathbf{q}=W_{0}.\end{split}$ (8)
+    The expansion coefficients are optimized to maximize the target performance metric according to (1) which is, using projections described in Step S.3, transformed into a solution to quadratically constrained quadratic program (QCQP) [^6] In later examples equality constraint is used, because the optimum is always reached with the equality.
+and reads
+    $\begin{split}\max\limits_{\mathbf{q}}\ &\mathbf{q}^{\mathrm{H}}\mathbf{A}\mathbf{q}+\mathrm{Re}{\mathbf{a}_{i}^{\mathrm{H}}\mathbf{q}}+\alpha,\\ \mathrm{s.t.}\ &\mathbf{q}^{\mathrm{H}}\mathbf{B}_{i}\mathbf{q}+\mathrm{Re}{\mathbf{b}_{i}^{\mathrm{H}}\mathbf{q}}+\beta\leq 0,\forall i\end{split}$ (5)
+    where one of the matrices $\mathbf{B}$ must be full rank and positive definite. This one is typically the total incident energy.
     The optimization task is resolved using standard methods of the QCQP.
 The optimization problem is convex, so the solution is unique [19], which ensures that the maximum is truly achieved [^7] The optimization problem is equivalent to that for determining performance limitations on antenna metrics in the frequency domain [29]..
-Within this paper, the optimization is more than a thousand times faster than the computation of the impulse response of the system.
+Within this paper, the optimization is more than a thousand times faster than the computation of the impulse response of the system (depending on discretization in spatial and frequency or domain).
 For more details about the convex optimization framework, see [30].
   5. S.5
-    The solution is found as optimal vector $\mathbf{q}$ in the frequency domain.
-The linear combination on the feed line then gives the optimal feeding vector which results in maximum peak radiated intensity in the frequency domain using the operators of step S.2.
-    The solution is transformed to the time domain via inverse Fourier transform (iFT) either as a single function or by weighted summation images of the basis functions $\xi_{n}(t)$.
-The far field is also transformed to depict the radiated pulse waveform in the time domain.
-Although described for far-field radiation scenarios, the methodology explained above is general and can easily be applied to different directions and extended to applications such as near-field energy focusing, wireless power transfer, or UWB radars.
-However, we intentionally omit design concerns related to these systems, such as topology optimization of the radiating and receiving antennas.
-We focus solely on feeding waveform optimization, which is, unlike topology optimization, a convex optimization problem.
-III-B Verification
-The steps mentioned above are applied to an example of a single dipole antenna.
+    The solution to (5) is found as optimal vector $\mathbf{q}$. The desired optimal excitation is then found as
+    $a_{p}(t)=\sum\limits_{n}q_{pn}\xi_{n}(t),$ (6)
+    where inverse Fourier transforms of basis functions $\xi_{n}(\omega)$ are known.
+  6. S.6
+    If the performance of the current system with the optimal excitation is not sufficient, a redesign is needed, since solely by excitation, the performance cannot be enhanced further.
+For that, automated computer-powered system optimization can be employed.
+Automated co-design of the excitation and system topology then allows for even better performance.
+    The automated co-design is computationally demanding and is realistic only when globally optimal excitations are found quickly. This is the case of a solution to QCQP developed in previous steps.
+IV Pulse Optimization Verification
+The steps mentioned above are applied to an example of a single dipole antenna and instanenous far-field intensity $\boldsymbol{F}(\boldsymbol{d},t)$ in direction $\boldsymbol{d}$ orthogonal to the dipole orientation.
+Since this serves as verification, the redesign step S.6 is omitted, and codesign loop is presented in Section VII.
 The calculation is simultaneously performed in the Antenna Toolbox for MATLAB [31] (AToM) and CST.
-The dipole with a length of $15\,\mathrm{mm}$ and a width of $3\,\mathrm{mm}$ is oriented along the $z$ axis.
+The dipole with a length of $L=150\,\mathrm{mm}$ and a width of $3\,\mathrm{mm}$ is oriented along the $z$ axis.
 The port is placed in the middle of the dipole and is connected to an ideal $50\,\mathrm{\Omega}$ feed line.
 Radiation in the spherical coordinate system is studied in the direction $\varphi=0$ and $\theta=\pi/2$.
+Far field consists of polarizations $\varphi$ and $\theta$.
 Azimuthal coordinate $\varphi$ references to angle from positive direction of $x$ axis in $xy$ plane and elevation coordinate $\theta$ references to angle from positive direction of $z$ axis.
-The antenna, including the port, is given, and all steps visualized in Fig.
+The antenna, including the port, is given with frequency range from $0\,\mathrm{GHz}$ to $3.8\,\mathrm{GHz}$ [25], and all steps visualized in Fig.
 1 can be performed.
-Choosing AToM and CST as the EM solvers and computing the far-field component $F_{\theta}$ in elevation direction for unity excitation at all frequencies results in Fig.
+Choosing AToM and CST as the EM solvers and computing the far-field along $\theta$ direction $F_{\theta}$ in elevation direction for unity excitation at all frequencies results in Fig.
 2.
-Figure 2: Far field of a dipole antenna with a length of $15\,\mathrm{mm}$ and width of $3\,\mathrm{mm}$ in directions $\varphi=0$ and $\theta=\pi/2$.
+Figure 2: Far field along $\theta$ direction $F_{\theta}$ of a dipole antenna with a length of $15\,\mathrm{mm}$ and width of $3\,\mathrm{mm}$ in directions $\varphi=0$ and $\theta=\pi/2$.
 The antenna is fed through a delta-gap source connected to the feed line, providing a voltage of $1\,\mathrm{V}$ across the entire frequency spectrum.
-Since the dipole is aligned along the $z$-axis, only the $\theta$ component of the electric field corresponds to radiation. The curves are in good agreement.
-In addition to the far field, an essential known characteristic is total incident energy, represented by the identity operator as seen in (4). This step accomplishes S.1.
+Since the dipole is aligned along the $z$-axis, only the $\theta$ component of the electric field corresponds to radiation and far-field along direction $\phi$ is vanishingly small.
+The curves are in good agreement.
+In addition to the far field, an essential characteristic is total incident energy. This step accomplishes S.1.
 Since the frequency data from CST are dense due to the 1001 samples, only linear interpolation is used.
-Data resulting from AToM are sparse and rational fitting is used for interpolation [^6] Vector fitting [41] could be applied to enforce passivity and causality..
-Metrics such as far field and total input power are projected to the band-limited functions defined in Appendix B resulting in matrices as denoted in S.3.
-The matrix operators allow for optimization S.4 which gives the optimal solution in the frequency domain.
-Since the basis functions used for projection are orthogonal and an identity operator represents total input power, the matrix for total input power is an identity matrix.
+Data resulting from AToM are sparse and rational fitting is used for interpolation [^8] Vector fitting [32] could be applied to enforce passivity and causality.
+according to S.2.
+Metrics such as far field and total input power are projected to the band-limited functions defined in Appendix C resulting in matrices as denoted in S.3,
+$\boldsymbol{F}(\boldsymbol{d},t)=\mathbf{F}\mathbf{q}.$ (7)
+The matrices allow for optimization S.4 which gives the optimal solution.
+Thanks to basis function orthonormality, the identity matrix of the incident energy is used as the full rank positive definite matrix in the QCQP,
+$\begin{split}\max\limits_{\mathbf{q}}\ &\mathbf{q}^{\mathrm{H}}\mathbf{F}^{\mathrm{H}}\mathbf{F}\mathbf{q},\\ \mathrm{s.t.}\ &\mathbf{q}^{\mathrm{H}}\mathbf{q}=W_{0}.\end{split}$ (8)
 The optimal excitation is a complex conjugate of the far field [1] in Fig.
 2.
-Total input energy is $W_{0}=10^{-10}\,\mathrm{J}$.
-Meeting the constraint in (8) is ensured by scaling the magnitude of the incident wave.
-The optimal waveform in the time domain is the result of iFT in S.5, and is plotted in Fig.3.
+For presentation purposes, the total input energy was set to $W_{0}=10^{-10}\,\mathrm{J}$.
+Using S.5, the optimal waveform in the time domain is plotted in Fig.3.
 Figure 3: Optimal waveform of the incident wave gained from analysis in different solvers. In both cases of AToM and CST, optimization is performed in MATLAB.
 The agreement of different solvers is more than satisfactory.
-The last step is the computation of the radiated pulse, shown in Fig. 4.
-Figure 4: Optimal radiated waveform for $t_{0}=0\,\mathrm{s}$ at distance $r$ gained from iFT of the far field for AToM data and monitoring the far field during simulation in CST, where the optimal incident wave is applied as excitation.
+The last step is the computation of the radiated pulse, shown as the magnitude of far-field along $\theta$ direction $F_{\theta}$ in Fig. 4.
+Figure 4: Far-field along $\theta$ direction $F_{\theta}$ optimal radiated waveform for $t_{0}=0\,\mathrm{s}$ at distance $r$ gained from iFT of the far field for AToM data and monitoring the far field during simulation in CST, where the optimal incident wave is applied as excitation.
 The similarity of the data computed by AToM with iFT and a synthetic measurement of the radiated pulse by optimal waveform excitation in CST verifies and, even more, generalizes the method for determining optimal pulses studied for decades [7, 9, 21, 14, 16].
 In addition, this framework allows us to find an optimal excitation using an arbitrary solver and do it for complex scenarios, as demonstrated in the following section.
-IV Representative Examples
-Multiple antennas and arrays are analyzed using the presented approach.
-Moreover, this section extends the study to antiferromagnetic memory switching where the target is the electric field magnitude at a chosen point excited by a pulsed incident field.
-Additional research is also conducted on the energy delivered to the given point during the chosen time interval.
-Since the solution is independent of the solver, as verified in the previous section, the most suitable one can be used for each scenario.
-IV-A Far-Field Pulsed Radiation
-IV-A1 Singleport Antennas
-As has been concluded in previous work, the bowtie dipole antenna exhibits good performance among planar antennas studied for pulsed radiation in the direction orthogonal to the design plane [26, 28].
+V Transmitting System
+A representative example of a transmitting system is a single port antenna radiating in the target direction in the far field.
+The methodology can be used to find the one with the best performance, i.e., with the highest allowed intensity of the instantaneous field.
+As has been concluded in previous work on pulsed farfield radiation, the bowtie dipole antenna exhibits good performance among planar antennas studied for pulsed radiation in the direction orthogonal to the design plane [14, 16].
 In this section, the following antennas are considered: a self-grounded bowtie antenna, a conical dipole antenna, a conical spiral antenna, and a logarithmic periodic antenna.
 The optimal waveform is determined to maximize the radiated peak power.
-None of the antennas are topologically optimized for the target.
+Incident power waves are used as degrees of freedom, which allows us to consider matching (reflected power waves).
+This is essential for a realistic pulse response.
 This example aims to demonstrate the utility of the approach and compare the best potentially achievable performance of the proposed antennas, including their optimal excitation waveform.
@@
 The antenna proposed in [34] is scaled to fit the circumscribing sphere mentioned above.
-The simulation is again performed in CST and shown in Fig.
+The simulation is again performed in CST and the antenna is shown in Fig.
 7.
@@
 The strip width equals $0.017$ times the length of the longest element and the number of elements is 20.
+The paremeters are manualy tuned to achieve better performance.
 The antenna is simulated using AToM and shown in Fig.
@@
 Figure 8: Logarithmic-periodic dipole wire antenna designed according to [35]. The antenna is discretized in the mesh of $1528$ triangles.
-The question remains as to which antenna performs better. A comparison is shown in Fig. 9.
-Figure 9: Normalized optimal radiated pointing vector at chosen distance $r$ of the compared antennas.
+The question remains as to which antenna performs better. A comparison of radiation intensity [36]
+$U(t)=\dfrac{1}{Z_{0}}\left|\boldsymbol{F}(t)\right|^{2}$ (9)
+is shown in Fig. 9.
+Figure 9: Optimal radiation intensity $U$ at chosen distance $r$ from the compared antennas.
 The logarithmic periodic antenna exceeds the other designs in performance.
@@
 This finding could prompt efforts to develop improved designs.
+This may be achieved through redesign, involving iterative tuning of the logarithmic-periodic dipole wire antenna.
+An advanced topology optimization way is shown in Section VII using an automated approach on a planar antenna.
 Regarding the origins of what constitutes good performance, the realized gain in the given direction accounts for the pulse’s sharpness and energy.
-The emerging differences of the antennas are depicted in Fig.
+The emerging differences in realized gain $G_{\mathrm{real}}$ of the antennas are depicted in Fig.
 10 which justifies the performance differences presented in Fig.
 9.
-Figure 10: The realized gain of the antenna, which incorporates both the antenna gain and the effects of impedance matching, is a critical factor for achieving optimal performance.
+Figure 10: The realized gain $G_{\mathrm{real}}$ of the antenna, which incorporates both the antenna gain and the effects of impedance matching, is a critical factor for achieving optimal performance.
 The highest and flattest realized gain gives the best performance which provides some direction for improved future designs.
+It is also noted that matching efficiency is a major part of the good realized gain of the logarithmic periodic antenna.
 In contrast, phase properties must be considered once an optimal excitation signal is generated.
@@
 However, the inclusion of these constraints may result in an increase in the duration of the optimization process.
-IV-A2 Multiport Systems
-The extension to multiport antennas or arrays is straightforward, as with dipole arrays within this section. The matrix formulation is also shown in (7), Appendices A and C.
-The goal of this section is to show whether end-fire or broad-side pulsed radiation perform better in the setup of the four active dipoles discussed in Section III.
-The dipoles are spaced half a wavelength at $1\,\mathrm{GHz}$.
-Again, a band limited from above by frequency $3.8\,\mathrm{GHz}$ is considered.
-The optimal radiated pulses of all configurations are compared in Fig.
-12 where the magnitudes of the radiated Poynting vector peaks $S$ are given and normalized by radial distance $r$.
-The results show that broadside radiation performs better than end-fire.
-Figure 12: Normalized Poynting vector of a single dipole antenna and array radiated in the end-fire and broadside direction.
-Figure 13: Optimal excitation signals for end-fire and broadside pulsed radiation.
-Considering the same input energy, the end-fire array increases the performance by more than a factor of four as compared to a single dipole.
-In the case of a broadside array, the performance boost reaches a factor of six.
-The difference may be due to the coupling of the field with the elements of the arrays which can be deduced from the optimal excitation waveforms in Fig.
-13.
-Although the end-fire and broadside arrays share the same configuration, their characteristics, especially impedance ratios, differ since the dipoles are fed simultaneously.
-In this particular case, the broadside direction performs better than the end-fire orientation.
-The excitations to all ports are almost identical in broadside operation.
-For the end-fire array, the delay equal to the distance between the dipoles multiplied by the speed of light is observed.
-The delays agree with the direction of end-fire radiation which points from dipole port 1 to dipole port 4.
-The magnitude of the excitation pulse is increased from dipole port 1 to dipole ports 2, 3, and 4, which indicates that the last element contributes the most.
-The suggested approach can be applied to multi-dimensional arrays or any radiating structures with multiple ports and arbitrary configurations.
-Typically, computational demands grow as the number of elements increases.
-As demonstrated, expanding the number of independent ports results in expressing the excitation waveforms at each port through the basis functions described in Appendix B.
-The expansion coefficients are distinctively grouped into a vector that includes all dof.
-IV-B Switching Antiferromagnetic Memory
-The approach mentioned in this manuscript can be generalized in an arbitrary application where the input-output relation is known.
-In this section, the switching of an antiferromagnetic memory using a device illuminated by a pulsed laser is analyzed.
+VI Receiving Systems
+Apart from transmission, the receiving performance is also covered by the approach.
+This shows that the method can also deal with the transmitting and receiving parts at the same time, but this is not shown in this manuscript.
+In this section, focused on terahertz pulses, the switching of an antiferromagnetic memory using a device illuminated by a pulsed laser, is analyzed.
 The data are taken from [11].
@@
 However, the simulation in CST is done only for the central part of the device, as seen in Fig.
-14, and the plane wave is considered as excitation.
+12, and the plane wave is considered as excitation.
 Due to the large width of the beam as compared to the target, the plane wave is an excellent approximation and simplifies the calculation.
@@
 The metric of interest is electric field intensity in CuMnAs layer $E_{\mathrm{CuMnAs}}$ in the center of the motive shown in Fig.
-14.
+12.
 The CuMnAs motive is prepared on a GaAs substrate and most of the CuMnAs alloy is covered by Au, except the central part as described in [11].
@@
 The excitation waveforms are plotted in Fig. 13.
-Figure 15: Incident plane waves at the position of the observation: (o) Pulse used in [45].
+Figure 13: Time course of incident plane waves $E_{\mathrm{inc}}$ at the position of the observation: (o) Pulse used in [11].
 (a) Optimal excitation for the maximum peak.
@@
 (d) Optimal excitation for the maximum peak with both constraints in (b) and (c).
-The pulses are normalized to the maximum value of $10^{5}\,\mathrm{V/cm}$. The excited field magnitude in CuMnAs is provided in Fig. 16.
-Figure 16: Magnitude of the electric field in CuMnAs for excitation pulses (o)–(d) in Fig. 15.
+The pulses are normalized to the maximum value of $10^{5}\,\mathrm{V/cm}$. The excited field magnitude in CuMnAs $E_{\mathrm{CuMnAs}}$ is provided in Fig. 14.
+Figure 14: Magnitude of the electric field in CuMnAs $E_{\mathrm{CuMnAs}}$ for excitation pulses (o)–(d) in Fig. 13.
 The pulse used in [11] is denoted as (o).
@@
 One of the optimal pulses presented corresponds to the maximum magnitude of the electric field in CuMnAs in $t=0$ which is denoted as (a).
-The optimization problem is identical to (8) other than matrix $\mathbf{S}$ being replaced by $\mathbf{E}^{\mathrm{H}}\mathbf{E}$, where the electric field at the target point is represented as $\boldsymbol{E}_{\mathrm{CuMnAs}}(\omega)=\mathbf{E}(\omega)\mathbf{q}$ in the frequency domain, and $\mathbf{q}$ corresponds to the expansion of incident plan wave $a(\omega)$ into the basis functions of Appendix B, as done in (5).
+The optimization problem is identical to (8), where $\mathbf{q}$ corresponds to the expansion of incident plane wave $a(\omega)$ into the basis functions of Appendix C, as done in (3).
 The relation is computed using CST which allows us to compute the electric field at a given point as a function of frequency considering unity excitation.
 Mapping the set of the given basis functions results in the desired matrix.
-The matrix representing the total energy of the incident plane wave is derived from (4) as demonstrated in (7), with the only variation being the inclusion of the free space impedance in the denominator.
+The matrix representing the total energy of the incident plane wave is derived in Appendix B.
 The maximum amplitude of the electric field in the CuMnAs alloy significantly exceeds that of the pulse used (o).
@@
 The incident wave is constrained to have $90\,\%$ energy in the interval $[-1.2,0.8]\,\mathrm{ps}$ taking into account the delay of $t_{\mathrm{d}}=0.2\,\mathrm{ps}$.
-The constraint is implemented using matrix $\mathbf{W}_{\mathrm{T}}$ of total available energy and energy available in the interval [2]
-$W_{\mathrm{T}}=\dfrac{1}{4\pi^{2}Z_{0}}\int\limits_{-\infty}^{\infty}\int\limits_{-\infty}^{\infty}\dfrac{2\sin\left[\left(\omega_{1}-\omega_{2}\right)T\right]}{\omega_{1}-\omega_{2}}\mathrm{e}^{-\mathrm{j}\left(\omega_{1}-\omega_{2}\right)t_{\mathrm{d}}}\\ a^{*}\left(\omega_{2}\right)a\left(\omega_{1}\right)\mathrm{d}\omega_{2}\mathrm{d}\omega_{1},$ (9)
-where $T=1\,\mathrm{ps}$ is the half-time interval and the asterisk represents a complex conjugation.
-In the frequency domain, the sinc function represents the process of time windowing.
-The matrix obtained is the outcome of the procedure illustrated in (7) regarding total input power.
+The constraint is implemented using matrix derived from operator given in Appendix B.
 The corresponding QCQP reads
@@
 $\begin{split}\max\limits_{\mathbf{q}}\ &\mathbf{q}^{\mathrm{H}}\mathbf{E}^{\mathrm{H}}\mathbf{E}\mathbf{q},\\ \mathrm{s.t.}\ &\mathbf{q}^{\mathrm{H}}\mathbf{W}\mathbf{q}=W_{0},\\ &\mathbf{q}^{\mathrm{H}}\left(0.9\mathbf{U}-\mathbf{U}_{\mathrm{T}}\right)\mathbf{q}\leq 0,\end{split}$ (11)
-where $\mathbf{U}$ and $\mathbf{U}_{\mathrm{T}}$ are matrices of the total squared electric field and the squared electric field in the given time interval for $T=7\,\mathrm{ps}$.
-The matrices are evaluated equivalently as $\mathbf{W}$ and $\mathbf{W}_{\mathrm{T}}$ except $t_{\mathrm{d}}$.
-The electric field at the target point as a function of frequency and unity excitation is taken from CST.
-The function is multiplied by its complex conjugate image and projected into the set of basis functions $\xi_{n}$ from Appendix B.
-$\mathbf{U}=\dfrac{1}{4\pi}\int\limits_{-\infty}^{\infty}a^{*}\left(\omega\right)\boldsymbol{E}_{\mathrm{CuMnAs}}^{\mathrm{H}}(\omega)\boldsymbol{E}_{\mathrm{CuMnAs}}(\omega)a\left(\omega\right)\mathrm{d}\omega.$ (12)
-Matrix $\mathbf{U}_{\mathrm{T}}$ differs only by introducing integration over two distinct angular frequencies $\omega_{1},\omega_{2}$ and the sinc function in the expression as in (9).
+where $\mathbf{U}$ and $\mathbf{U}_{\mathrm{T}}$ are matrices of the total squared electric field and the squared electric field in the given time interval $T=7\,\mathrm{ps}$ similarly derived as delivered energy matrices in Appendix B.
 The results are presented in (c) which shows the excitation again spanning a long time, whereas the waveform in CuMnAs outside of the prescribed interval is attenuated compared to the amplitude global maxima which are not far from the optimum without the constraint.
@@
 This showcases the effectiveness of the tool in configuring pulses for antiferromagnetic memory switching and judging topologies for their capabilities.
-V Conclusion
+VII Optimization of the System’s Topology
+Since the excitation waveform optimization is fast and robust, it can be combined with the topology optimization.
+The target application is antenna and far-field radiation.
+The antenna is a multiport system, and this example, as a by-product, proves its applicability to systems with multiple inputs [^9] Multiple output systems within the objective can be treated by multiobjective optimization [30], whose solutions represent Pareto optimal points..
+The initial example consists of an array of two parallel and identical dipoles with the same dimensions as in Section IV.
+The frequency band also remains the same.
+The goal is to maximize instantaneous radiation in the end-fire direction.
+The simplicity of the design enables rapid assessment of the impulse response and facilitates the transfer of data between iterations without requiring re-evaluation in a full-wave solver, which is beneficial when integrating system and waveform optimization.
+Application to larger structures with multiple ports does not bring any complications except for computation time.
+By performing a parametric sweep, the optimal spacing between the dipoles is determined. Different configurations are evaluated:
+  • Both dipoles are driven.
+  • One dipole is driven, the other is shorted and acts as a reflector.
+  • One dipole is driven, the other is shorted and acts as a director.
+  • One dipole is driven and the reflector is, in its geometrical center, loaded by an optimized reactive loading.
+The maximum instantaneous radiation intensity achieved by these setups is shown in Fig.
+15 as a function of the distance between the dipoles, which is normalized to twice the dipole length, corresponding to the wavelength for which the dipole is designed.
+Figure 15: Comparison of maximum instantaneous radiation intensity $U$ for an end-fire two-dipole array for different distances between the dipoles normalized by the length of the dipoles.
+Performance is benchmarked against a single dipole.
+Considered configurations include both elements being driven, one element being driven with either a shorted director or reflector, and a reflector with an optimized load or a reflector setup with optimized topology.
+In the top panel, the solid lines show performance without considering impedance mismatch, reactive loading, or shape optimization, while, in the bottom panel, the dashed lines account for practical matching to an ideal $50\,\Omega$ feeding line, providing a more realistic assessment.
+The dotted line in the bottom panel represents a dipole reflector with an optimized reactive load, and the markers in the bottom panel represent reflector setups with optimized shapes (a)-(d) in Fig.
+16.
+Both optimizations in the bottom panel consider connection to ideal $50\,\Omega$ feed line.
+[…]
+[…]
+[…]
+Figure 16: Topologically optimized antenna shapes with ports highlighted in green. Instantaneous radiation intensity in the north direction is optimized (reflector setup):
+Shape (a) Best performance among the presented designs.
+Shape (b) Elements of the reflector closest to the driven element.
+Shape (c) Good trade-off between the shape regularity and performance.
+Shape (d) The simplest and the most compact design.
+Figure 15 shows two panels:
+  • In the top panel, the solid lines with better performance omit the matching (assume ideal voltage generators at all frequencies).
+The dof are delta-gap voltages.
+In that case, the instantaneous far-field intensity is normalized to the total dissipated energy due to ohmic losses in the elements and radiation.
+  • In the bottom panel, the dashed lines with lower performance consider the reflection of waves at the ports i.e., matching.
+The dof are incident power waves in the $50\,\Omega$ feeding lines.
+There, the instantaneous far-field intensity is normalized to the total incident energy.
+As expected, the best performance is achieved when both dipoles are fed independently.
+A similar trend is followed by a single fed dipole with a shorted reflector; for larger distances, the impact of the reflector is negligible, and the performance is identical to that of a dipole alone.
+The director decreases performance as compared to a single dipole.
+Continuing in the co-design, the performance of the setup with reflector can be improved by loading the reflector with a shorted ideal $50\,\Omega$ line of an optimized length $\ell$ in range from $0\,\mathrm{m}$ to $2L$ (reactive loading).
+The optimization of reactive loading is only considered in the realistic scenario, where reflection caused by mismatch is taken into account, and the optimized dof is the incident power wave.
+The enhancement is significant at small separations between the driven dipole and the reflector.
+The optimized design performance is plotted as a dotted line in Fig.
+15 and can be compared to the dashed line for the shorted reflector.
+The best possible performance is $36.2\,\mathrm{mW}$ which corresponds to $d/(2L)=0.11$ and line length $\ell/(2L)=0.20$.
+In the same way, the load on the director was also optimized, but the performance did not exceed that of a single dipole, so the result is not depicted.
+The ultimate solution is co-design with the topology optimization [37, 38], which is implemented for the reflector setup.
+Performances are shown as markers in Fig.
+15 for designs (a) – (d) from Fig.
+16, respectively.
+The antennas possess east-west mirror symmetry with respect to the north target direction.
+Designs are ranked according to their performance, inversely related to their regularity [39].
+Performance enhancement is significant and underscores the effectiveness of the co-design methodology.
+This approach effectively addresses mismatch issues, improves gain, and exceeds the capabilities of reactive loading in terms of functionality.
+VIII Conclusion
 A methodology to determine optimal waveforms is presented in this paper with optimized parameters being defined as linear or quadratic terms.
 The evaluation requires a post-processing step based on convex optimization and the input data can be calculated in any EM simulator.
-The tool can serve and provide quantitative results in many scenarios, including pulsed radiation or antiferromagnetic memory switching with pulse lasers, as discussed in this paper, as well as the creation of hotspots using near-field energy focusing in lossy materials using antenna arrays for applications such as hyperthermia therapy in medicine.
+The technique can serve and provide quantitative results in many scenarios, including pulsed radiation or antiferromagnetic memory switching with pulse lasers, as discussed in this paper, as well as the creation of hotspots using near-field energy focusing in lossy materials using antenna arrays for applications such as hyperthermia therapy in medicine.
 The capability of the tool is to find the optimal excitation waveform for an arbitrarily chosen target metric.
@@
 By decoupling waveform optimization from structural design, the proposed method ensures a globally optimal solution that can be computed efficiently in polynomial time, unlike the iterative and often local solutions associated with topology optimization.
-This characteristic guarantees the general applicability of the method to existing antenna systems without necessitating their redesign, while still ensuring the best achievable performance for a given scenario and set of constraints.
-While this paper focuses on waveform optimization for a fixed transmitting antenna, the methodology has broader potential.
-It can be extended to co-design scenarios in which both antenna and waveform design are considered simultaneously, making it applicable to a wider range of advanced engineering problems.
-In such case, the topology optimization would change the design according to the results computed by the proposed tool, though such a procedure would be computationally intensive.
-However, this would open the way to the automated design of pulsed radiators, antiferromagnetic memories, or other devices that rely on pulses in the time domain.
+This characteristic guarantees the general applicability of the method to existing systems without necessitating their redesign, while still ensuring the best achievable performance for a given scenario and set of constraints.
+The methodology has broader potential.
+It is extended to co-design scenarios in which both antenna and waveform design are considered simultaneously, making it applicable to a wider range of advanced engineering problems.
+In such case, the topology optimization changes the design according to the results computed by the proposed tool.
+The combination of the pulse and topology design is present on an end-fire two-dipole array, resulting in a significant improvement of the performance.
 Appendix A Matrix Operators from EFIE and MoM
 A representative example of an EM solver capable of providing required antenna characteristics is AToM based on EFIE and MoM.
-The matrix representation of multiple quantities is presented in [39, Appendix A].
+The matrix representation of multiple quantities is presented in [30, Appendix A].
 The extraction of essential metrics to the manuscript is presented in this appendix.
@@
 where $c$ is the speed of light, $\mathbf{F}$ is a far-field matrix, and $\mathbf{I}$ is a column vector of current density expansion coefficients.
-The essential quantities are far field as a function of the incident waves from the feeding transmission line and their total available energy in (4).
+The essential quantities are far field as a function of the incident waves from the feeding transmission line and their total available energy.
 The electric far field is computed from far-field matrix $\mathbf{F}$ and current vector $\mathbf{I}$, see (12).
@@
 The study of the characteristics provides information on the suitability of individual antennas for application.
-Appendix B Band Limited Set of Basis Functions
+Appendix B Energy and Field Operators
+Fourier transform is used to transform target quantities from the time domain to the frequency domain. The mapping applies to energy and field expressions.
+The field operator for a given location or direction, and time reads
+$\boldsymbol{E}(\boldsymbol{r},t)=\\ \int\limits_{-\infty}^{\infty}\mathbf{H}\left(\boldsymbol{r},\omega\right)\mathbf{a}\left(\omega\right)\mathrm{e}^{\mathrm{j}\omega t}\,\mathrm{d}\omega.$ (15)
+One of the common targets is intensity $\left|\boldsymbol{E}\left(\boldsymbol{r},t\right)\right|^{2}$ at optimized time $t_{\mathrm{opt}}$.
+Choosing an arbitrary point in the space $\boldsymbol{r}$ and time window of length $2T$, another important measure is energy delivered to that point in that time window which reads
+$W(\boldsymbol{r},T)=C\int\limits_{t_{\mathrm{d}}-T}^{t_{\mathrm{d}}+T}\left|\boldsymbol{E}\left(\boldsymbol{r},t\right)\right|^{2}\mathrm{d}t=\\ C\int\limits_{\Omega}\int\limits_{\Omega}\dfrac{2\sin\left[\left(\omega_{1}-\omega_{2}\right)T\right]}{\omega_{1}-\omega_{2}}\mathrm{e}^{-\mathrm{j}\left(\omega_{1}-\omega_{2}\right)t_{\mathrm{d}}(\boldsymbol{r})}\\ \mathbf{a}^{\mathrm{H}}\left(\omega_{2}\right)\mathbf{H}^{\mathrm{H}}(\boldsymbol{r},\omega_{2})\mathbf{H}(\boldsymbol{r},\omega_{1})\mathbf{a}\left(\omega_{1}\right)\mathrm{d}\omega_{2}\mathrm{d}\omega_{1},$ (16)
+where $\Omega:\omega\in[-\omega_{\mathrm{max}},-\omega_{\mathrm{min}}]\cup[\omega_{\mathrm{min}},\omega_{\mathrm{max}}]$ is the intgration domain, $C$ is a constant to fit the units and $t_{\mathrm{d}}$ is the time delay from the excitation source.
+For incident energy, the transfer function $\mathbf{H}$ is omitted and the delay time is zero
+$W_{\mathrm{in}}(\boldsymbol{r},T)=C\int\limits_{t_{\mathrm{d}}-T}^{t_{\mathrm{d}}+T}\left|\mathbf{a}\left(t\right)\right|^{2}\mathrm{d}t=\\ C\int\limits_{\Omega}\int\limits_{\Omega}\dfrac{2\sin\left[\left(\omega_{1}-\omega_{2}\right)T\right]}{\omega_{1}-\omega_{2}}\mathbf{a}^{\mathrm{H}}\left(\omega_{2}\right)\mathbf{a}\left(\omega_{1}\right)\mathrm{d}\omega_{2}\mathrm{d}\omega_{1}.$ (17)
+The expression is simplified for the total energies, which read,
+$W(\boldsymbol{r},T)=C\int\limits_{-\infty}^{\infty}\left|\boldsymbol{E}\left(\boldsymbol{r},t\right)\right|^{2}\mathrm{d}t=\\ C\int\limits_{\Omega}\mathrm{e}^{-\mathrm{j}\omega t_{\mathrm{d}}(\boldsymbol{r})}\mathbf{a}^{\mathrm{H}}\left(\omega\right)\mathbf{H}^{\mathrm{H}}(\boldsymbol{r},\omega)\mathbf{H}(\boldsymbol{r},\omega)\mathbf{a}\left(\omega\right)\mathrm{d}\omega,$ (18)
+or
+$W_{\mathrm{in}}(\boldsymbol{r},T)=C\int\limits_{-\infty}^{\infty}\left|\mathbf{a}\left(t\right)\right|^{2}\mathrm{d}t=C\int\limits_{\Omega}\mathbf{a}^{\mathrm{H}}\left(\omega\right)\mathbf{a}\left(\omega\right)\mathrm{d}\omega,$ (19)
+Galerkin method [27, 28] converts continuous operators shown above to matrices using the same basis and testing functions.
+The method is applied to all operators using functions introduced in Appendix C.
+Appendix C Band Limited Set of Basis Functions
 This manuscript uses a set of basis functions for the spectral expansion of antenna characteristics.
@@
 The values $\omega_{\mathrm{min}}$ and $\omega_{\mathrm{max}}$ define the spectral band interval where the functions are nonzero.
-Appendix C Projection via Galerkin Method
-Galerkin method [38, 13] converts continuous operators to matrices using the same basis and testing functions.
-The method is applied to all operators using functions introduced in Appendix B.
-The approach is explicitly presented on the far-field operator and the construction of the far-field power matrix.
-The far-filed operator is projected in the following way for optimized time $t_{\mathrm{opt}}$
-$\hat{\mathbf{F}}_{pn}(\boldsymbol{r},\omega)=\\ \int\limits_{-\infty}^{\infty}\mathbf{F}\left(\boldsymbol{r},\omega\right)\mathbf{Z}^{-1}\left(\omega\right)\mathbf{P}\mathbf{k}_{\mathrm{i}}^{-1}\left(\omega\right)\xi_{n}\left(\omega\right)\mathrm{e}^{\mathrm{j}\omega t_{\mathrm{opt}}}\,\mathrm{d}\omega,$ (17)
-where $\mathbf{a}_{p}$ excites only the chosen feeding transmission line $p$ by unity amplitude.
-Here, it is presented for operators from MoM, see Appendix A, but the same methodology can be applied to antenna characteristics from an arbitrary EM solver.
-The far-field operator projected has two rows corresponding to polarizations and $pn$ columns with a linear index that combines ports and basis functions.
-The far-filed operator allows us to define the far-field power matrix
-$\mathbf{S}(\boldsymbol{r},\omega)=\hat{\mathbf{F}}^{\mathrm{H}}(\boldsymbol{r},\omega)\hat{\mathbf{F}}(\boldsymbol{r},\omega)$ (18)
-in (8).